### v1

# Bimanual Deformable Bag Manipulation Using a Structure-of-Interest Based Latent Dynamics Model

Abstract
The manipulation of deformable objects by robotic systems presents a significant challenge due to their complex and infinite-dimensional configuration spaces. This paper introduces a novel approach to Deformable Object Manipulation (DOM) by emphasizing the identification and manipulation of Structures of Interest (SOIs) in deformable fabric bags. We propose a bimanual manipulation framework that leverages a Graph Neural Network (GNN)-based latent dynamics model to succinctly represent and predict the behavior of these SOIs. Our approach involves constructing a graph representation from partial point cloud data of the object and learning the latent dynamics model that effectively captures the essential deformations of the fabric bag within a reduced computational space. By integrating this latent dynamics model with Model Predictive Control (MPC), we empower robotic manipulators to perform precise and stable manipulation tasks focused on the SOIs. We have validated our framework through various empirical experiments demonstrating its efficacy in bimanual manipulation of fabric bags. Our contributions not only address the complexities inherent in DOM but also provide new perspectives and methodologies for enhancing robotic interactions with deformable objects by concentrating on their critical structural elements. Experimental videos can be obtained from https://sites.google.com/view/bagbot
Index Terms:
Deformable object manipulation, structure of interest, latent dynamics model, bimanual manipulation.
I Introduction
Deformable object manipulation (DOM) [1, 2, 3] is a fundamental capability for robots to meaningfully interact with the physical world and assist in various human tasks. However, the manipulation of deformable objects such as cloth [4], rope [5], and food ingredients[6] is particularly challenging due to their infinite-dimensional configuration space and complex dynamics. Traditional methods in DOM often resort to simplified physics models or data-driven modeling with handcrafted features, which lack adaptability for the varied shapes and dynamics of these objects [7]. Moreover, most current DOM works focus on manipulating the entire object, neglecting the critical structures, i.e., Structures of Interest (SOI), that are essential for subsequent manipulation steps.
Figure 1: (Left) SOI Examples for different deformable object manipulation tasks, e.g., garment hanging, robot-assistive dressing. (Right) Conceptual representation of the manifold with boundary. The manifold encompasses $\operatorname{Int}\mathcal{M}$ and $\partial\mathcal{M}$, where the local neighborhoods of points in $\operatorname{Int}\mathcal{M}$ and ${\partial\mathcal{M}}$ are homeomorphically equivalent to $\operatorname{Int}\mathbb{H}^{n}$ and $\partial\mathbb{H}^{n}$.
Figure 2: The two robots grasp two handles of a fabric bag to manipulate the SOI (i.e., the opening rim) into the target configuration.
In this paper, we introduce the concept of SOI into the realm of DOM (see Fig. 1 for an example), a paradigm shift that emphasizes the importance of identifying and manipulating key structural components rather than the entire object. This focus on the SOI is motivated by the observation that successful DOM tasks typically involve the manipulation of these key areas. By targeting these SOIs, we can reduce the computational load significantly, as modeling the complete 3D dynamics of the deformable object is unnecessary and burdensome for the task at hand.
Furthermore, this work is pioneering in considering the bimanual manipulation of a deformable fabric bag as shown in Fig. 2. By proposing a novel bimanual manipulation framework using a GNN-based latent dynamics model, we address the complexities of DOM with a focus on SOIs. Our approach is designed to extract the SOI from the observed object point cloud, construct a graph representation, and learn the latent dynamics model that effectively captures the object’s deformations within a compact space. Integrating this model with model predictive control (MPC), we enable robots to achieve accurate and stable manipulation of deformable bags.
The main contributions of this work include:
  • Pioneering the bimanual manipulation of deformable fabric bags by focusing on SOIs.
  • Introducing the SOI concept for representing deformable object states, which emphasizes the manipulation of critical structures.
  • Designing a GNN-based method to learn a latent dynamics model from partial point cloud data, particularly focusing on the SOIs.
  • Implementing MPC based on the latent dynamics for generating optimal manipulation actions centered around SOIs.
Various empirical experiments on the bimanual manipulation of a fabric bag validate the efficacy of our proposed framework. Our work offers new insights and methodologies for improving robotic systems’ capability for intelligent deformable object manipulation, with a specialized emphasis on the pivotal SOIs.
Figure 3: Bimanual bag manipulation is formulated as a POMDP problem, where the SOI-related observation $\hat{o}_{t}$ is extracted from the original observation $o_{t}$ and governed by $f_{\theta_{\mathrm{dym}}}$.
II Related Work
Deformable object manipulation has been an active research area in robotics. Existing methods can be categorized into model-based and data-driven approaches. Model-based methods rely on simplified physics models to represent deformable objects. Early works used mass-spring models (MSM) to simulate deformation [8, 9]. The finite element method (FEM) provides more accurate modeling of continuum mechanics [10, 11, 12]. However, analytical models require extensive manual tuning and generalization across different materials or shapes remains difficult. Data-driven methods [7, 13, 14] aim to learn models directly from data. Vision-based methods extract geometric features from visual observations to infer deformations [15, 16]. Recent works utilized deep learning on point cloud data and achieved improved modeling accuracy [17]. However, they depend heavily on large labeled datasets. Self-supervised methods were proposed to learn from physical interactions [18, 19]. However, they focused on planar objects and could not handle complex deformations.
Our work is mostly related to robot manipulation using graph neural networks (GNNs). GNNs have shown promising results in learning physics simulations [20, 21, 22, 23, 24] and control policies [25, 26, 27, 28]. Recently, GNNs were introduced to model rope and cloth manipulation [29, 30, 31]. Different from these works, we propose the structure of interest concept to focus on key object components and employ GNNs to learn latent dynamics models for describing complex deformable object behaviors. The integration of the latent dynamics model with MPC also distinguishes our framework from prior arts. In summary, our work aims to advance existing deformable object manipulation by introducing an SOI-based modeling approach. The adoption of state-of-the-art deep learning techniques allows better generalizability across different materials and shapes. The experiments on the bimanual manipulation of fabric bags represent challenging test cases and validate the feasibility of our framework.
III Problem Statement
Given that individual image and depth observations generally do not fully disclose the state of the environment, we approach the task of bimanual bag manipulation as a Partially Observable Markov Decision Process (POMDP) as depicted in Fig. 3. This is formally defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{O},\mathcal{B},\mathcal{R},\gamma)$, where the state at time $t$, denoted by $s_{t}$, belongs to the state space $\mathcal{S}$ and is not directly observable. The state encapsulates the configuration of the robots and the manipulated object. The corresponding observation at time $t$, denoted by $o_{t}$, is within the observation space $\mathcal{O}$. The state transition model $\mathcal{T}(s_{t+1}\mid s_{t},a_{t})$ describes the probability of transitioning from the current state $s_{t}$ to a new state $s_{t+1}$ upon taking an action $a_{t}$ from the action space $\mathcal{A}$, which consists of the combined left and right robotic actions, represented by the Cartesian product $\mathcal{A}=\mathcal{A}^{1}\times\mathcal{A}^{2}$. The function $\mathcal{B}(o_{t}\mid s_{t},a_{t-1})$ specifies the likelihood of observing $o_{t}$ after executing action $a_{t-1}$ and transitioning to state $s_{t}$. The reward function $\mathcal{R}(s_{t},a_{t})$ assigns a valued reward to each state-action pair, and the discount factor $\gamma\in[0,1)$ quantifies the preference for immediate rewards over future rewards.
The goal of this work is to use two 3D-printed robot grippers to grasp the handles to manipulate the opening rim to achieve a target state $\mathbf{g}$. We assume this deformable bag manipulation task is a quasi-static manipulation and dynamic manipulation motions are not considered. At time step $t$, the robots apply action $(\mathbf{a}^{1}_{t},\mathbf{a}^{2}_{t})\in\mathcal{A}$ upon the bag, and we can partially observe transitions of the bag from $\mathbf{o}_{t}$ to $\mathbf{o}_{t+1}$ under the unknown state transitions from $\mathbf{s}_{t}$ to $\mathbf{s}_{t+1}$. However, a complete observation of the bag is not necessary, in our task, the opening rim of the bag is critical for successful manipulation tasks since it not only determines the manipulation task goals but also provides the most informative sensory feedback, such as visual landmarks, during manipulation. We define the Structure of Interest (SOI) in the context of Deformable Object Manipulation (DOM) refers to specific regions or features of a deformable object that are critical for successful manipulation tasks (see Fig. 1 for examples.). Therefore, in this task, we consider the opening rim of the manipulated bag as our SOI points, and topologically, we can define this loop-like structure as a manifold with boundary. As illustrated in Fig. 1, we also define its interior and boundary as $\operatorname{Int}\mathcal{M}$ and ${\partial\mathcal{M}}$, whose points’ neighborhoods are respectively homeomorphic to $\operatorname{Int}\mathbb{H}^{n}=\{(x_{1},\ldots,x_{n})\mid x_{n}>0\}$ and $\partial\mathbb{H}^{n}=\{(x_{1},\ldots,x_{n})\mid x_{n}=0\}$.
With an appropriate perception module, the observation of the Structure of Interest (SOI), denoted by $\hat{o}_{t}$, can be extracted from the overall observation of the bag $o_{t}$. Our approach is based on the insight that it is more efficacious to predict the dynamics of the SOI rather than the entire complex dynamics of the bag. To this end, we employ a Graph Neural Network (GNN) to establish a dynamics model $f_{\theta_{\text{dym}}}$ that is dedicated to learning the transition functions of the SOI, defined as $f_{\theta_{\text{dym}}}:\hat{\mathcal{O}}\times\mathcal{A}\rightarrow\hat{\mathcal{O}}$. This dynamics model accepts as input a sequence of SOI observations $\mathbf{\hat{o}}_{t-n,\ldots,t}\in\mathcal{\hat{O}}$ and actions $(\mathbf{a}^{1}_{t-n,\ldots,t},\mathbf{a}^{2}_{t-n,\ldots,t})\in\mathcal{A}$, and predicts the subsequent observation $\mathbf{\hat{o}}_{t+1}$, where $n$ represents the length of the observation history before the current time step $t$. With the dynamics model, we proceed to cast the bimanual manipulation of the bag as a task within the Model Predictive Control (MPC) framework. Within this MPC setup, the cost function $\mathcal{J}$ quantifies the difference between the final SOI feature points at time step $T$ and the targeted SOI state $\mathbf{g}$. Details on the precise structure of the cost function $\mathcal{J}$ are illustrated in the following section. This cost is minimized to yield an optimal sequence of actions across a temporal horizon of $T$ steps:
$\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{0:{T-1}}$ $=\underset{\langle\mathbf{a}^{0},\mathbf{a}^{1}\rangle_{0:T-1}\in\mathcal{A}}{\arg\min}\mathcal{J}(\mathbf{\hat{g}},\mathbf{g})$ (1)
$\text{where}\quad\mathbf{\hat{g}}$ $=f_{\theta_{dym}}\left(\mathbf{\hat{o}}_{0},\langle\mathbf{a}^{0},\mathbf{a}^{1}\rangle_{0:{T-1}}\right)$
where $\hat{\mathbf{g}}$ represents the predicted SOI state at time step $T$. Fig. 4 shows the overall framework of our proposed latent dynamics model for the bimanual bag manipulation task.
Figure 4: The conceptual representation of the proposed framework for bimanual deformable fabric bag manipulation is based on the latent SOI dynamics model.
Figure 5: The proposed SOI global particle sampling process. Left: the reconstructed point cloud for the fabric bag. Right: (a) raw SOI point cloud. (b) preprocessed SOI point cloud. (c) Reconstructed SOI Surface. (d) Resampled SOI particles.
IV Methodology
IV-A Global Particle Sampling from Raw Observation
In this work, we begin by sampling representative particles from raw visual inputs, i.e., RGB-D images, to facilitate the training of Graph Neural Network (GNN)-based dynamics model, depicted in Fig. 4(a). The challenge lies in extracting meaningful particle representations of the Structure of Interest (SOI) from visual data that is significantly occluded. To this end, we introduce a comprehensive global particle sampling methodology, as illustrated in Fig. 5.
Preprocessing: The initial phase of our methodology involves refining the raw point cloud data obtained from RealSense D435 cameras. With appropriate calibration, we can transform the raw RGBD images into point cloud data with the cameras’ intrinsic and extrinsic parameters. Through meticulous calibration, we can convert raw RGB-D images into point cloud data, utilizing the intrinsic and extrinsic parameters of the cameras. To enhance data quality, a statistical outlier removal filter is applied to strip away noise, paving the way for accurate analysis. Focusing on the bag’s opening rim, which is crafted from green fabric, we employ a color-based segmentation algorithm. A specific color threshold within the HSV color space is set to discern green hues, effectively isolating the rim from the rest of the bag. Concurrently, we identify and omit the points marking the bag’s handle areas where robots will interact with the fabric using a separate HSV threshold. To ensure the rim’s point cloud is distinctly segregated from other bag components, Euclidean clustering is performed as a concluding step, particularly if handle detection leaves any connected segments. This meticulous process results in a purified dataset, optimal for the rim’s geometric representation.
Surface Reconstruction: To reconstruct the green rim’s surface from its point cloud, we leverage Poisson reconstruction, a preferred technique for its proficiency in managing occlusions by interpolating over incomplete data to form a smooth, continuous surface. Traditional methods like ball pivoting and alpha shapes underperform in occluded scenarios. Before reconstruction, we compute point normal via local principal axis analysis to guide the Poisson algorithm. The resulting surface is further polished with a uniform resampling algorithm, ensuring mesh uniformity and feature preservation for subsequent modeling and manipulation tasks.
Refinement with Topological Prior: In this step, we ensure that the reconstructed surface conforms to the anticipated topological structure of the bag’s opening rim through a two-fold approach: 1) Topological Analysis: We incorporate topological priors by recognizing that the rim should form a contiguous loop. This knowledge is employed to scrutinize the mesh, detecting any topological discrepancies. 2) Corrective Actions: Upon identifying any topological deviations, we undertake corrective measures by cutting and remeshing to realign the mesh with a loop topology that accurately represents the bag’s rim.
Resampling: In the final stage of our process, our objective is to generate a high-quality particle set that accurately represents the geometry of the rim. To achieve this, we initially implement voxel grid down-sampling to attain a uniform distribution of vertices while concurrently eliminating statistical outliers. Subsequently, we employ the farthest point sampling technique to selectively reduce the point cloud associated with the Structure of Interest (SOI) to a manageable quantity conducive to Graph Neural Network (GNN) training. The culmination of this process, resulting in a refined particle dataset, is depicted in Fig. 5(d).
IV-B SOI Dynamics Model
To characterize the dynamics of SOI particles, as illustrated in Fig. 4(b), we begin to construct a particle graph and introduce the graph neural networks to model the SOI dynamics. To construct a particle graph, we denote a graph representing SOI observations as $\mathcal{G}(\mathbf{\hat{o}}_{t})=(V_{t},E_{t})$, and the graph vertices $V_{t}$ correspond to the SOI’s particles $\mathbf{p}_{i,t}$. Each particle is expressed as $\mathbf{p}_{i,t}=\langle\mathbf{x}_{i,t},\mathbf{b}_{i,t}\rangle$, where the position and attributes of the particle $i$ at time $t$ are denoted as $\mathbf{x}_{i,t}$ and $\mathbf{b}_{i,t}$, respectively. Edges $E_{t}$ link vertices dynamically, based on spatial relationships, connecting all neighboring particles within a predefined range. Edge relations are captured by $\mathbf{e}_{j}=\langle m_{j},n_{j},\mathbf{c}_{j}\rangle$, with $1\leq m_{j},n_{j}\leq|\mathbf{P}_{t}|$ denoting the indices of the connected particles, $j$ being the edge index, and $\mathbf{c}_{j}$ describing the type of connection, be it internal structural or handle-to-rim linkages.
The goal of introducing GNNs is to simulate the dynamics of the SOI and to predict subsequent states from a short historical sequence of SOI particle graphs, formalized as:
$\mathcal{G}(\mathbf{\hat{o}}_{t+1})=f_{\theta_{\text{dyn}}}\left(\mathcal{G}(\mathbf{\hat{o}}_{t-h:t}),\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{t}\right)$ (2)
To facilitate this, the original high-dimensional observation space concerning the SOI is reduced into a condensed, low-dimensional latent graph by encoding the distinct particle and connection features of the SOI. The loss function of each encoder is defined based on a distance metric as:
$\phi^{\mathbf{p}},\psi^{\mathbf{p}}$ $=\arg\min_{\phi^{\mathbf{v}},\psi^{\mathbf{v}}}\operatorname{Dist}[V-(\phi^{\mathbf{p}}\circ\psi^{\mathbf{p}})(P)]$ (3)
$\phi^{\mathbf{e}},\psi^{\mathbf{e}}$ $=\arg\min_{\phi^{\mathbf{e}},\psi^{\mathbf{e}}}\operatorname{Dist}[E-(\phi^{\mathbf{e}}\circ\psi^{\mathbf{e}})(E)]$
Here, $\phi^{\mathbf{p}}:V\rightarrow Z_{V}$ and $\phi^{\mathbf{e}}:E\rightarrow Z_{E}$ serve as the encoders for SOI particles and edges, while $\psi^{\mathbf{p}}:Z_{V}\rightarrow V$ and $\psi^{\mathbf{e}}:Z_{E}\rightarrow E$ function as their respective decoders. The encoding of SOI particle and edge features via $\phi^{\mathbf{p}}$ and $\phi^{\mathbf{e}}$ is executed as follows:
$z_{i,t}^{\mathbf{p}}$ $=\phi^{\mathbf{p}}(\mathbf{p}_{i,t})$ (4)
$z_{j,t}^{\mathbf{e}}$ $=\phi^{\mathbf{e}}(\mathbf{p}_{m_{j},t},\mathbf{p}_{n_{j},t},\mathbf{c}_{j})$
Subsequently, the dynamics are captured using the decoders $\psi^{\mathbf{p}}$ and $\psi^{\mathbf{e}}$, leading to the prediction of the SOI particle graph at time $t+1$:
$\mathbf{\hat{e}}_{k,t}$ $=\psi^{\mathbf{e}}(z_{j,t}^{\mathbf{e}})_{j=1,\cdots,|E_{t}|}$ (5)
$\hat{\mathbf{p}}_{i,t+1}$ $=\psi^{\mathbf{p}}\left(z_{i,t}^{\mathbf{p}},\sum_{j\in\mathcal{N}_{i}}\mathbf{\hat{e}}_{k,t}\right)_{i=1,\cdots,|\mathbf{P}_{t}|}$
Within this context, $\mathcal{N}_{i}$ denotes the set of edges where particle $i$ is the recipient. To adeptly handle the instantaneous propagation of forces, the training also integrates multistep message passing.
Our training data, derived from point clouds, lacks a consistent point-to-point mapping across each frame, thus precluding the use of particle-wise loss functions. To quantify the resemblance between two sets of SOI particle distributions, we investigate two distinct loss functions. The first is the widely-adopted Chamfer distance (CD), computed between two particle sets $\mathbf{P}_{1},\mathbf{P}_{2}\subseteq\mathbb{R}^{3}$ as follows:
$\small\mathcal{L}_{\text{CD}}(\mathbf{P}_{1},\mathbf{P}_{2})=\sum_{\mathbf{x}_{1}\in\mathbf{P}_{1}}\min_{\mathbf{x}_{2}\in\mathbf{P}_{2}}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{2}^{2}+\sum_{\mathbf{x}_{2}\in\mathbf{P}_{2}}\min_{\mathbf{x}_{1}\in\mathbf{P}_{1}}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{2}^{2}$ (6)
The second is the Earth mover’s distance (EMD), which is formulated as:
$\small\mathcal{L}_{\text{EMD}}(\mathbf{P}_{1},\mathbf{P}_{2})=\min_{\phi:\mathbf{P}_{1}\rightarrow\mathbf{P}_{2}}\sum_{\mathbf{x}_{1}\in\mathbf{P}_{1}}\|\mathbf{x}_{1}-\phi(\mathbf{x}_{1})\|_{2}$ (7)
where $\phi:\mathbf{P}_{1}\rightarrow\mathbf{P}_{2}$ denotes a bijective mapping. The EMD represents a solution to the assignment problem, ensuring a unique and stable optimal bijection $\phi$ for almost every pair of particle sets, invariant to infinitesimally small point displacements. In essence, EMD in our framework aligns distributions while mitigating point cloud anomalies through the definition of bijective correspondence. Regarding the Chamfer distance, it’s worth noting that its use here is somewhat liberal, as it does not fulfill the triangle inequality property. Our composite loss function integrates these distances in a weighted fashion: $\mathcal{L}(\mathbf{P}_{1},\mathbf{P}_{2})=\alpha\mathcal{L}_{\text{CD}}(\mathbf{P}_{1},\mathbf{P}_{2})+\beta\mathcal{L}_{\text{EMD}}(\mathbf{P}_{1},\mathbf{P}_{2})$. Empirical evaluations suggest that the optimal weights are $\alpha=0.85$ and $\beta=0.15$.
IV-C Model Predictive Control
Upon training our SOI-centric latent dynamics model, we integrate a model predictive control (MPC) approach to control the robotic gripper in manipulating the fabric bag, as depicted in Fig. 4(c). We simplify the gripper’s action space into a parameterized form: $(x,y,z,r_{z})$, with $\{x,y,z\}$ representing the end-effector’s position interacting with the bag handles, and $r_{z}$ indicating the gripper’s rotation around the vertical $z$ axis. We omit $r_{x}$ and $r_{y}$ rotations based on empirical evidence suggesting minimal SOI deformation from these movements. A goal-oriented MPC is employed as below, using $\mathbf{g}$ to symbolize the desired SOI shape and $\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{0:{T-1}}$ to represent the sequence of action pairs derived via MPC, with $T$ as the planning horizon.
$\min_{\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{0:T-1}}$ $\mathcal{J}\left(\mathbf{\hat{o}}_{T},\mathbf{g}\right)=\min_{\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{0:T-1}}\mathcal{L}\left(\mathbf{\hat{P}}_{T},\mathbf{P}_{\mathbf{g}}\right)$ (8)
s.t. $\mathcal{G}(\mathbf{\hat{o}}_{t+1})=f_{\theta_{\text{dym}}}\left(\mathcal{G}(\mathbf{\hat{o}}_{t-h:t}),\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{t}\right)$
$\mathbf{P}_{\mathbf{g}}=\mathcal{G}(\mathbf{g})$
$\mathbf{\hat{P}}_{t}=\mathcal{G}(\mathbf{\hat{o}}_{t})$
$t=0,1,\ldots,T-1$
The objective is to identify action sequences that minimize the hybrid distance to the goal, expressed as $\mathcal{J}(\mathbf{\hat{g}}_{T},\mathbf{g})\equiv\mathcal{L}(\mathbf{\hat{P}}_{T},\mathbf{P}_{\mathbf{g}})$. Gradient-based trajectory optimization is utilized to ascertain the minimal-cost trajectory, showcased in Fig. 4(c). Initial random shooting in the simplified action space is followed by cost calculation via the GNN-driven dynamics model. Subsequently, gradient refinement with the limited-memory BFGS method [32] is conducted on the most cost-effective trajectories, employing the same loss function used during the dynamics model’s training phase.
Figure 6: Experimental set-up. (a) Overview of the experimental set-up to validate our SOI-based dynamics model for bimanual deformable bag manipulation tasks. (b)-(d) The SOI shape categories are considered in experiments: Long Oval (LO), Round Oval (RO), and Short Oval (SO).
TABLE I: Averaged SOI particle sampling results across 90 frames.
Sampling Methods CD(cm) $\downarrow$ EMD(cm) $\downarrow$ GD(cm) $\downarrow$
Global Sampling (GPS) $1.68\pm 0.21$ $1.27\pm 0.35$ $2.14\pm 0.32$
Local Sampling (LPS) $\mathbf{1.46}\pm 0.11$ $\mathbf{1.08}\pm 0.22$ $\mathbf{1.97}\pm 0.23$
V Experiments
V-A Experiment Setup
As shown in Fig. 6, we present the general experimental setup used to validate our SOI-based dynamics model for bimanual deformable bag manipulation tasks. The setup includes four RealSense D435 RGBD cameras, each positioned at a corner, to capture RGBD images of the deformable fabric bag from multiple angles at a frequency of 30Hz and a resolution of 640$\times$480. Two Dobot CR5 robotic manipulators, equipped with 7 degrees of freedom, grasp the bag’s handles using 3D-printed grippers secured with zip ties. The Structure of Interest (SOI) of the bag, specifically the opening rim, has been modified for better SOI perception: it is cut and sewn with the same type of fabric but in a contrasting green color, which facilitates the detection while maintaining the consistency of the bag’s dynamic behavior. Additionally, a top-down camera and a front-facing camera are employed to capture the experimental process from various perspectives.
A dataset of 18,000 frames, around 30 minutes, comprises the training data, featuring 200 episodes, each spanning 90 frames. Within every episode, we execute six actions on the fabric bag. Before each episode’s inception, the fabric bag is orderly arranged to ensure the opening rim protrudes outward without excessive wrinkling. The data collection policy indiscriminately selects from the action space parameters after doing collision checking and bag constraint checking, which include translations along the $x$, $y$, $z$ axes, and the rotation around $rz$. Throughout each episode, we archive partial point clouds captured by RGB-D cameras, as well as the robot end-effector’s poses as returned by the robot controller. Data recording is contingent on the gripper’s movement, an approach taken to optimize memory consumption.
As illustrated in Fig. 6(b)-(d), this study considers three SOI shape categories: Long Oval (LO), Round Oval (RO), and Short Oval (SO). The experiment comprises two types of manipulation tasks. In the SOI shape preservation task, the robotic manipulators aim to keep the SOI shape consistent while moving the fabric bag from its initial position to a predefined target location. Both the initial and target configurations share the same shape but differ in their spatial transformations. Conversely, the SOI shape servoing task involves altering the SOI shape from one configuration to another distinct shape configuration by manipulating the bag handles. We conduct a quantitative evaluation of each component within our framework, encompassing SOI particle sampling, the SOI dynamics model, and performance results of designed manipulation tasks. Our primary metrics for evaluating the manipulation performance are Chamfer distance (CD) and Earth mover’s distance (EMD), along with a hybrid metric that combines both. Furthermore, we employ Geodesic distance (GD) [20] to assess the proximity of boundary points residing on one manifold with boundary to those on another.
TABLE II: The performance of the SOI particle dynamics model with different loss function.
Loss Functions CD(cm) $\downarrow$ EMD(cm) $\downarrow$ GD(cm) $\downarrow$
CD $1.52\pm 0.19$ $1.94\pm 0.18$ $2.53\pm 0.25$
EMD $1.85\pm 0.17$ $\mathbf{1.37}\pm 0.15$ $2.76\pm 0.13$
$0.2$ CD + $0.8$ EMD $1.59\pm 0.16$ $1.40\pm 0.17$ $2.13\pm 0.18$
$0.1$ CD + $0.9$ EMD $1.55\pm 0.17$ $1.43\pm 0.16$ $2.28\pm 0.19$
$0.15$ CD + $0.85$ EMD $\mathbf{1.50}\pm 0.16$ $1.38\pm 0.17$ $\mathbf{2.02}\pm 0.16$
Figure 7: Qualitative results in (a) SOI shape preserving and (b) SOI shape servoing experiments.
V-B SOI Particle Sampling
We commence by benchmarking the proposed Global Particle Sampling (GPS) technique against the Local Particle Sampling (LPS) baseline. LPS initiates by processing the SOI-specific partial point cloud of the fabric bag through its unique color signature. Subsequently, it encapsulates the incomplete SOI-specific cloud with a convex hull. Following this, it performs the point sampling procedure, eventually amalgamating the sampled points into a full set of SOI-specific particles. As shown in Table I, we compute the mean distance between the sampled particles and the ground-truth particles, the latter captured by a professional 3D scanner. Our analysis indicates that the GPS method incurs lower losses in terms of all distance metrics, thus outperforming the LPS approach. These outcomes align with the premise that incorporating additional topological information about the SOI can markedly enhance the quality of sampling, particularly in scenarios where occlusions are present.
V-C GNN-based SOI Dynamics Model
The training of our GNN model, detailed in Section IV-B, commences with the construction of a graph, where edges are formed between vertices that are within a distance threshold of $d=0.04$. Every vertex and edge are encoded using 3-layer Multilayer Perceptrons (MLPs), featuring hidden and output layers, each with 300 neurons. The propagation module is constituted by a fully connected neural layer with a layer size of 300. For motion prediction, we employ an additional 3-layer MLP with the hidden layer configured to 300 neurons. ReLU activation functions are utilized throughout the neural networks to introduce non-linearity. The model undergoes training for 120 epochs, employing the Adam optimizer. We have selected a batch size of 32 to balance the trade-off between generalization and computational efficiency. The learning rate is set at $5\mathrm{e}$-4, which is a conventional choice for steady convergence. These hyperparameters were chosen to foster a robust learning process while maintaining the capacity to capture complex patterns within the data.
Subsequently, we evaluate the performance of GNN-based SOI dynamics models using different loss functions. Table 6 indicates that only using CD or EMD individually optimizes for its own metric but not others, but combining CD and EMD losses leads to better overall performance on all metrics, and the Geodesic distance is significantly improved, compared to using them individually. The combination of $0.85~{}\text{CD}+0.15~{}\text{EMD}$ achieves the best performance across all metrics, giving the lowest CD, competitive EMD, and lowest GD, so we select this hybrid distance for the following experiments. The reason is that CD and EMD losses capture different aspects of point cloud similarity. CD compares point-wise distances while EMD measures global shape differences. By combining them, the model is optimized for both local point accuracy and global shape matching. The proper combination of 0.85 CD + 0.15 EMD balances these objectives best.
TABLE III: Mean Hybrid Distance Error for SOI Shape Preserving by Different Dynamics Modeling Methods
Method Long Oval (LO) Round Oval (RO) Short Oval (SO)
MSM $2.63\pm 0.31$ $2.82\pm 0.38$ $2.94\pm 0.41$
FEM $1.96\pm 0.27$ $2.13\pm 0.36$ $2.42\pm 0.38$
Ours $\textbf{1.69}\pm 0.18$ $\textbf{1.84}\pm 0.22$ $\textbf{2.08}\pm 0.21$
TABLE IV: Mean Hybrid Distance Error and Success Rate for SOI Shape Servoing by Different Methods
Method Mean Hybrid Distance Error Success Rate
LO $\rightarrow$ SO SO $\rightarrow$ RO LO $\rightarrow$ RO LO $\rightarrow$ SO SO $\rightarrow$ RO LO $\rightarrow$ RO Total
VS [33] $3.75\pm 0.19$ $3.69\pm 0.22$ $3.92\pm 0.23$ $21/30$ $19/30$ $22/30$ $68.89\%$
LSC [34] $2.94\pm 0.48$ $3.14\pm 0.42$ $2.89\pm 0.44$ $24/30$ $24/30$ $26/30$ $82.23\%$
MSM [35] $3.22\pm 0.43$ $3.49\pm 0.47$ $3.28\pm 0.39$ $19/30$ $21/30$ $23/30$ $70.00\%$
FEM [36] $2.61\pm 0.36$ $2.82\pm 0.39$ $2.67\pm 0.33$ $23/30$ $22/30$ $26/30$ $78.89\%$
Ours $2.25\pm 0.23$ $2.39\pm 0.28$ $2.13\pm 0.24$ $28/30$ $/30$ 30 $29/30$ 96.67%
V-D Manipulation Results
To evaluate the performance of the proposed SOI-based dynamics model for deformable fabric bag manipulation, we conducted SOI preserving and servoing experiments on a fabric bag with different oval shapes and the transitions between them. In SOI preserving experiments, we mainly compare the performance of our proposed approach with two commonly-used and well-established dynamics modeling techniques: the Mass-Spring Model (MSM) [35] and the Finite Element Model (FEM) [36]. The comparative analysis encompasses both quantitative and qualitative outcomes, as encapsulated in Figure 7(a) and Table III. Our findings reveal a discernible trend across all examined methods; there is an incremental rise in error corresponding to the transformation of the bag’s shape from a Long Oval (LO) to a Short Oval (SO), suggesting that as the bag opening rim becomes shorter and wider, the complexity of preserving its structure increases. Notably, our GNN-based dynamics model consistently outperforms the MSM and FEM across the entire spectrum of object shapes, achieving the lowest error rates in tasks dedicated to shape preservation. This outcome corroborates the superior expressive capability of our GNN-based model in capturing the intricate deformations of the fabric bags, and maintaining the SOI shape unchanged during the fabric bag moving, surpassing the traditional MSM and FEM approaches, particularly when dealing with objects that exhibit more complex dynamic behaviors. The advantage of our model is most pronounced when interacting with simpler geometric shapes.
In our SOI servoing experiments, we conduct a comprehensive analysis of the manipulation results associated with transitions between distinct shape categories of the structure of interest (SOI). Additionally, we juxtapose our dynamics model with two sophisticated shape servoing techniques tailored for deformable objects–visual servoing (VS) as described by Lagneau et al. [33], and the latent shape control (LSC) model introduced by Qi et al. [34]. The analysis integrates both quantitative and qualitative assessments, as depicted in Fig. 7(b) and detailed in Table IV. We consider three specific shape servoing tasks: LO (Long Oval) $\rightarrow$ SO (Short Oval), SO $\rightarrow$ RO (Round Oval), and LO $\rightarrow$ RO. The evaluation metrics include the Mean Hybrid Distance Error (recorded solely for successful trials) and the Success Rate of SOI servoing. Furthermore, we report the success rate for achieving complete servoing with a threshold error of less than $5$ cm, based on 30 trials for each method and shape transition scenario. The experimental outcomes reveal that our proposed method consistently secured the least shape error, ranging from $2.13$ to $2.39$ cm across all tested cases, and achieved an outstanding overall success rate of $96.67\%$. When contrasted with traditional model-based techniques such as the Mass-Spring Model (MSM) and the Finite Element Model (FEM), our data-driven SOI dynamics model exhibited a markedly enhanced capability in manipulating the SOI for the fabric bag. It was observed that transformations involving shorter and wider shapes incurred higher errors due to their inherently complex dynamics. Nevertheless, our proposed method demonstrated a robust performance across all tasks. In conclusion, the bimanual manipulation experiments underscore the effectiveness of the proposed SOI-based dynamics modeling approach, which ensures precise and dependable control over the shapes of deformable objects through advanced predictive modeling and optimization techniques.
VI Conclusion
In this work, we introduced a novel bimanual manipulation framework for deformable bags, centered around a Structure of Interest (SOI)-based latent dynamics model. Our approach effectively integrates multi-view perception, graph neural networks, and model predictive control, facilitating precise and efficient robotic manipulation of flexible materials. Through simulations and real-world experiments, the framework demonstrated promising results in intelligent physical interaction with deformable objects. One limitation of the current system is its reliance on differently colored SOIs to distinguish target manipulation areas, which may not be feasible in all operational settings. Future efforts will aim to enhance the system’s ability to identify and manipulate SOIs without such visual aids, broadening the applicability of our method to a wider array of real-world scenarios.

### v2

@@
-Bimanual Deformable Bag Manipulation Using a Structure-of-Interest Based Latent Dynamics Model
+Bimanual Deformable Bag Manipulation Using a Structure-of-Interest Based Neural Dynamics Model
 Abstract
-The manipulation of deformable objects by robotic systems presents a significant challenge due to their complex and infinite-dimensional configuration spaces.
-This paper introduces a novel approach to Deformable Object Manipulation (DOM) by emphasizing the identification and manipulation of Structures of Interest (SOIs) in deformable fabric bags.
-We propose a bimanual manipulation framework that leverages a Graph Neural Network (GNN)-based latent dynamics model to succinctly represent and predict the behavior of these SOIs.
-Our approach involves constructing a graph representation from partial point cloud data of the object and learning the latent dynamics model that effectively captures the essential deformations of the fabric bag within a reduced computational space.
-By integrating this latent dynamics model with Model Predictive Control (MPC), we empower robotic manipulators to perform precise and stable manipulation tasks focused on the SOIs.
-We have validated our framework through various empirical experiments demonstrating its efficacy in bimanual manipulation of fabric bags.
-Our contributions not only address the complexities inherent in DOM but also provide new perspectives and methodologies for enhancing robotic interactions with deformable objects by concentrating on their critical structural elements.
-Experimental videos can be obtained from https://sites.google.com/view/bagbot
+The manipulation of deformable objects by robotic systems presents significant challenges due to their complex dynamics and infinite-dimensional configuration spaces.
+This paper introduces a novel approach to deformable object manipulation (DOM) by emphasizing the introduction and manipulation of structures of interest (SOIs) in deformable fabric bags.
+We propose a bimanual manipulation framework that leverages a graph neural network (GNN)-based neural dynamics model to succinctly represent and predict the behavior of these SOIs.
+Our approach involves global particle sampling process to construct a particle representation from partial point clouds of the SOIs and learning the neural dynamics model that effectively captures the essential deformations of the SOIs for fabric bags.
+By integrating this neural dynamics model with model predictive control (MPC), we enable robotic manipulators to perform precise and stable manipulation tasks focused on the SOIs.
+We validate our new framework through various experiments that demonstrate its efficacy in manipulating deformable bags and T-shirts.
+Our contributions not only address the complexities inherent in DOM but also provide new perspectives and methodologies for enhancing robotic interactions with deformable materials by concentrating on their critical structural elements.
+Videos of the conducted experiments can be seen at https://sites.google.com/view/bagbot.
 Index Terms:
-Deformable object manipulation, structure of interest, latent dynamics model, bimanual manipulation.
+Deformable object manipulation, structure of interest, neural dynamics model, bimanual manipulation.
 I Introduction
 Deformable object manipulation (DOM) [1, 2, 3] is a fundamental capability for robots to meaningfully interact with the physical world and assist in various human tasks.
-However, the manipulation of deformable objects such as cloth [4], rope [5], and food ingredients[6] is particularly challenging due to their infinite-dimensional configuration space and complex dynamics.
-Traditional methods in DOM often resort to simplified physics models or data-driven modeling with handcrafted features, which lack adaptability for the varied shapes and dynamics of these objects [7].
-Moreover, most current DOM works focus on manipulating the entire object, neglecting the critical structures, i.e., Structures of Interest (SOI), that are essential for subsequent manipulation steps.
-Figure 1: (Left) SOI Examples for different deformable object manipulation tasks, e.g., garment hanging, robot-assistive dressing.
+However, the manipulation of deformable objects such as cloths [4], ropes [5], and food materials [6] is particularly challenging due to their infinite-dimensional configuration space and complex dynamics.
+Traditional methods in DOM often resort to simplified physics models or data-driven modeling with handcrafted features, which lack adaptability for the varied shapes and dynamics of these objects.
+Moreover, most current DOM works focus on manipulating the entire object, neglecting the critical structures, i.e., structures of interest (SOIs), that are essential for subsequent manipulation steps.
+Figure 1: (Left) SOI Examples for different deformable object manipulation tasks, e.g., garment hanging, and robot-assistive dressing.
 (Right) Conceptual representation of the manifold with boundary.
@@
 In this paper, we introduce the concept of SOI into the realm of DOM (see Fig.
-1 for an example), a paradigm shift that emphasizes the importance of identifying and manipulating key structural components rather than the entire object.
+1 for an example), a paradigm shift that emphasizes the importance of leveraging and manipulating key structural components rather than the entire object.
 This focus on the SOI is motivated by the observation that successful DOM tasks typically involve the manipulation of these key areas.
 By targeting these SOIs, we can reduce the computational load significantly, as modeling the complete 3D dynamics of the deformable object is unnecessary and burdensome for the task at hand.
-Furthermore, this work is pioneering in considering the bimanual manipulation of a deformable fabric bag as shown in Fig.
+In this work, we address the bimanual manipulation of a deformable fabric bag as illustrated in Fig.
 2.
-By proposing a novel bimanual manipulation framework using a GNN-based latent dynamics model, we address the complexities of DOM with a focus on SOIs.
-Our approach is designed to extract the SOI from the observed object point cloud, construct a graph representation, and learn the latent dynamics model that effectively captures the object’s deformations within a compact space.
-Integrating this model with model predictive control (MPC), we enable robots to achieve accurate and stable manipulation of deformable bags.
-The main contributions of this work include:
-  • Pioneering the bimanual manipulation of deformable fabric bags by focusing on SOIs.
-  • Introducing the SOI concept for representing deformable object states, which emphasizes the manipulation of critical structures.
-  • Designing a GNN-based method to learn a latent dynamics model from partial point cloud data, particularly focusing on the SOIs.
-  • Implementing MPC based on the latent dynamics for generating optimal manipulation actions centered around SOIs.
-Various empirical experiments on the bimanual manipulation of a fabric bag validate the efficacy of our proposed framework.
-Our work offers new insights and methodologies for improving robotic systems’ capability for intelligent deformable object manipulation, with a specialized emphasis on the pivotal SOIs.
-Figure 3: Bimanual bag manipulation is formulated as a POMDP problem, where the SOI-related observation $\hat{o}_{t}$ is extracted from the original observation $o_{t}$ and governed by $f_{\theta_{\mathrm{dym}}}$.
+Our proposed framework introduces a GNN-based neural dynamics model that enables to tackle the challenges associated with deformable object manipulation (DOM), with an emphasis on Structures of Interest (SOIs).
+The framework seeks to effectively represent the state of the object by extracting SOIs’ particles from the point cloud data, and learning the underlying dynamics within a condensed particle space.
+By integrating this method with model predictive control (MPC), we enable robots to achieve accurate and stable manipulation of deformable bags.
+The contributions of this work are as follows:
+  • The development of an original bimanual manipulation strategy for deformable fabric bags with an emphasis on the utility of SOIs.
+  • The adoption of the SOI concept for deformable object representations using the proposed global particle sampling method, highlighting the significance of key structural elements during manipulation.
+  • A new approach to learning the SOI-based neural dynamics models via GNNs, applied to particle sets of SOIs.
+  • The application of MPC informed by the neural dynamics model to guide the generation of optimal manipulation actions that prioritize SOIs.
+Our experiments demonstrate the potential of this framework in the context of bimanual manipulating deformable objects, such as fabric bags, T-shirts.
+While our findings suggest improvements in the robotic manipulation of deformable objects by concentrating on SOIs, we acknowledge the ongoing need for research in this area.
+This work contributes to the broader understanding of intelligent manipulation strategies and opens pathways for future innovations in the robotic handling of complex materials.
 II Related Work
-Deformable object manipulation has been an active research area in robotics.
+Deformable object manipulation (DOM) has been an active research area in robotics.
 Existing methods can be categorized into model-based and data-driven approaches.
 Model-based methods rely on simplified physics models to represent deformable objects.
-Early works used mass-spring models (MSM) to simulate deformation [8, 9].
-The finite element method (FEM) provides more accurate modeling of continuum mechanics [10, 11, 12].
+Early works used mass-spring models (MSM) to simulate deformation [7].
+The finite element method (FEM) provides more accurate modeling of continuum mechanics [8, 9].
 However, analytical models require extensive manual tuning and generalization across different materials or shapes remains difficult.
-Data-driven methods [7, 13, 14] aim to learn models directly from data.
+Data-driven methods [10, 11] aim to learn models directly from data.
 Vision-based methods extract geometric features from visual observations to infer deformations [12, 13].
 Recent works utilized deep learning on point cloud data and achieved improved modeling accuracy [14].
-However, they depend heavily on large labeled datasets.
+However, they depend heavily on large labelled datasets.
 Self-supervised methods were proposed to learn from physical interactions [15, 16].
 However, they focused on planar objects and could not handle complex deformations.
-Our work is mostly related to robot manipulation using graph neural networks (GNNs).
-GNNs have shown promising results in learning physics simulations [20, 21, 22, 23, 24] and control policies [25, 26, 27, 28].
-Recently, GNNs were introduced to model rope and cloth manipulation [29, 30, 31].
-Different from these works, we propose the structure of interest concept to focus on key object components and employ GNNs to learn latent dynamics models for describing complex deformable object behaviors.
-The integration of the latent dynamics model with MPC also distinguishes our framework from prior arts.
-In summary, our work aims to advance existing deformable object manipulation by introducing an SOI-based modeling approach.
-The adoption of state-of-the-art deep learning techniques allows better generalizability across different materials and shapes.
-The experiments on the bimanual manipulation of fabric bags represent challenging test cases and validate the feasibility of our framework.
+Recent advancements in DOM have leveraged graph neural networks (GNNs) to model the complex interactions within varying materials and shapes [17, 18, 19, 20].
+Wang et al.
+[21] proposed an offline-online learning framework that utilizes GNNs for the deformation model in cable manipulation.
+This method captures the intricate physics of cables, achieving a balance between offline learning from a rich dataset and online refinement through interaction.
+Deng et al.
+[22] introduced a deep reinforcement learning approach that employs local GNNs for a goal-conditioned rearranging of deformable objects.
+Their method efficiently adapts to the changing dynamics of the objects during manipulation tasks.
+In the domain of elasto-plastic object manipulation, Shi et al.
+[23] presented RoboCraft, which combines perception and simulation through GNNs to understand and shape 3D objects.
+Furthermore, the GDOOM framework by Ma et al.
+[24] focuses on learning latent graph dynamics for visual manipulation, offering a comprehensive approach to predicting deformable object behavior.
+In the specific area of deformable bag manipulation, a number of approaches have been proposed.
+Xu et al.
+[10] developed DextAIRity, a system that simplifies the manipulation of deformable objects by harnessing pneumatic controls.
+Gu et al.
+[25] introduced ShakingBot, which demonstrates dynamic manipulation skills for bagging tasks, highlighting the potential to handle high-speed and complex deformable object interactions.
+Chen et al.
+[26] focused on the challenge of opening plastic bags and inserting objects, a task with significant practical applications in automation.
+Lastly, Weng et al.
+[27] explored the concept of interactive perception for deformable object manipulation, where the perception and interaction are tightly integrated to enhance manipulation capabilities.
+Our work distinguishes itself from the existing literature by introducing a novel bimanual manipulation framework for deformable fabric bags that places a strong emphasis on structures of interest (SOIs).
+Unlike previous work that focused on the manipulation of entire objects or specific applications such as fabric or plastic bag handling, our approach centers on identifying and manipulating the SOIs within fabric bags to achieve precise and efficient DOM.
+We propose a GNN-based latent dynamics model tailored to represent and predict the behavior of these SOIs, which is a subtle yet meaningful advancement from the general-purpose models in the current literature.
+Furthermore, by integrating this model with model predictive control (MPC), our framework enables accurate and stable manipulation of fabric bags by robotic systems in a computationally efficient manner.
+Our contributions lie in the specialized focus on SOIs and the combination of GNN-based modeling with MPC, subtly enriching the landscape with fresh viewpoints for enhancing robotic DOMs.
+Figure 3: (Left) Conceptual representation of SOI-based bimanual deformable object manipulation problem; (Right) The problem is formulated as a POMDP setting, where the SOI-related observation $\mathbf{o}^{soi}_{t}$ is extracted from the original observation $o_{t}$ and governed by $f_{\theta_{\mathrm{dyn}}}$.
 III Problem Statement
@@
 3.
-This is formally defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{O},\mathcal{B},\mathcal{R},\gamma)$, where the state at time $t$, denoted by $s_{t}$, belongs to the state space $\mathcal{S}$ and is not directly observable.
-The state encapsulates the configuration of the robots and the manipulated object.
+This is formally defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{O},\Omega,\mathcal{R},\gamma)$, where the state at time $t$, denoted by $s_{t}$, belongs to the state space $\mathcal{S}$ and is not directly observable.
+The state encapsulates the configuration of the robots and the manipulated deformable object.
 The corresponding observation at time $t$, denoted by $o_{t}$, is within the observation space $\mathcal{O}$.
-The state transition model $\mathcal{T}(s_{t+1}\mid s_{t},a_{t})$ describes the probability of transitioning from the current state $s_{t}$ to a new state $s_{t+1}$ upon taking an action $a_{t}$ from the action space $\mathcal{A}$, which consists of the combined left and right robotic actions, represented by the Cartesian product $\mathcal{A}=\mathcal{A}^{1}\times\mathcal{A}^{2}$.
-The function $\mathcal{B}(o_{t}\mid s_{t},a_{t-1})$ specifies the likelihood of observing $o_{t}$ after executing action $a_{t-1}$ and transitioning to state $s_{t}$.
+The state transition model $\mathcal{T}(s_{t+1}\mid s_{t},a_{t})$ describes the probability of transitioning from the current state $s_{t}$ to a new state $s_{t+1}$ upon taking an action $a_{t}$ from the action space $\mathcal{A}$, which consists of the combined left and right robotic actions, represented by the Cartesian product $\mathcal{A}=\mathcal{A}^{0}\times\mathcal{A}^{1}$.
+The function $\Omega(o_{t}\mid s_{t},a_{t-1})$ specifies the likelihood of observing $o_{t}$ after executing action $a_{t-1}$ and transitioning to state $s_{t}$.
 The reward function $\mathcal{R}(s_{t},a_{t})$ assigns a valued reward to each state-action pair, and the discount factor $\gamma\in[0,1)$ quantifies the preference for immediate rewards over future rewards.
-The goal of this work is to use two 3D-printed robot grippers to grasp the handles to manipulate the opening rim to achieve a target state $\mathbf{g}$.
-We assume this deformable bag manipulation task is a quasi-static manipulation and dynamic manipulation motions are not considered.
-At time step $t$, the robots apply action $(\mathbf{a}^{1}_{t},\mathbf{a}^{2}_{t})\in\mathcal{A}$ upon the bag, and we can partially observe transitions of the bag from $\mathbf{o}_{t}$ to $\mathbf{o}_{t+1}$ under the unknown state transitions from $\mathbf{s}_{t}$ to $\mathbf{s}_{t+1}$.
+The goal of this work is to use two robot manipulators to firmly grasp and manipulate the SOI of the deformable object to achieve a target SOI state $\mathbf{y}_{*}$.
+We assume this deformable object manipulation task is a quasi-static manipulation and dynamic manipulation motions are not considered.
+As shown in Fig.
+3, at time step $t$, the dual-arm manipulators apply action $\{\mathbf{a}^{0}_{t},\mathbf{a}^{1}_{t}\}=\mathbf{a}_{t}\in\mathcal{A}$ upon the deformable object, and we can partially observe transitions of the object from $\mathbf{o}_{t}$ to $\mathbf{o}_{t+1}$ under the unknown state transitions from $\mathbf{s}_{t}$ to $\mathbf{s}_{t+1}$.
 However, a complete observation of the bag is not necessary, in our task, the opening rim of the bag is critical for successful manipulation tasks since it not only determines the manipulation task goals but also provides the most informative sensory feedback, such as visual landmarks, during manipulation.
-We define the Structure of Interest (SOI) in the context of Deformable Object Manipulation (DOM) refers to specific regions or features of a deformable object that are critical for successful manipulation tasks (see Fig.
-1 for examples.).
-Therefore, in this task, we consider the opening rim of the manipulated bag as our SOI points, and topologically, we can define this loop-like structure as a manifold with boundary.
+We define the structure of Interest (SOI) in the context of deformable object manipulation (DOM) refers to specific regions or features of a deformable object that are critical for successful manipulation tasks (e.g., examples in Fig.
+1).
+Therefore, in this task, we consider the opening rim of the manipulated bag as our SOI points, and topologically, we can define this loop-like structure as a manifold with boundary [lee2012smooth].
 As illustrated in Fig.
 1, we also define its interior and boundary as $\operatorname{Int}\mathcal{M}$ and ${\partial\mathcal{M}}$, whose points’ neighborhoods are respectively homeomorphic to $\operatorname{Int}\mathbb{H}^{n}=\{(x_{1},\ldots,x_{n})\mid x_{n}>0\}$ and $\partial\mathbb{H}^{n}=\{(x_{1},\ldots,x_{n})\mid x_{n}=0\}$.
-With an appropriate perception module, the observation of the Structure of Interest (SOI), denoted by $\hat{o}_{t}$, can be extracted from the overall observation of the bag $o_{t}$.
-Our approach is based on the insight that it is more efficacious to predict the dynamics of the SOI rather than the entire complex dynamics of the bag.
-To this end, we employ a Graph Neural Network (GNN) to establish a dynamics model $f_{\theta_{\text{dym}}}$ that is dedicated to learning the transition functions of the SOI, defined as $f_{\theta_{\text{dym}}}:\hat{\mathcal{O}}\times\mathcal{A}\rightarrow\hat{\mathcal{O}}$.
-This dynamics model accepts as input a sequence of SOI observations $\mathbf{\hat{o}}_{t-n,\ldots,t}\in\mathcal{\hat{O}}$ and actions $(\mathbf{a}^{1}_{t-n,\ldots,t},\mathbf{a}^{2}_{t-n,\ldots,t})\in\mathcal{A}$, and predicts the subsequent observation $\mathbf{\hat{o}}_{t+1}$, where $n$ represents the length of the observation history before the current time step $t$.
+With an appropriate perception module, the observation of the SOI, denoted by $\mathbf{o}^{soi}_{t}$, can be extracted from the overall observation of the bag $\mathbf{o}_{t}$.
+Our approach is based on the insight that it is more efficient to predict the dynamics of the SOI rather than the entire complex dynamics of the bag.
+To this end, we employ a graph neural network (GNN) to establish a dynamics model $f_{\theta_{\text{dyn}}}$ that is dedicated to learning the transition functions of the SOI, defined as $f_{\theta_{\text{dyn}}}:\mathcal{O}^{soi}\times\mathcal{A}\rightarrow\mathcal{O}^{soi}$.
+This dynamics model accepts as input a sequence of SOI observations $\mathbf{o}^{soi}_{t-n:t}\in\mathcal{O}^{soi}$ and bimanual actions $\{\mathbf{a}^{0},\mathbf{a}^{1}\}_{t-n:t}\subseteq\mathcal{A}$, and predicts the subsequent SOI observation $\mathbf{o}^{soi}_{t+1}$, where $n$ represents the length of the observation history before the current time step $t$.
 With the dynamics model, we proceed to cast the bimanual manipulation of the bag as a task within the model predictive control (MPC) framework.
-Within this MPC setup, the cost function $\mathcal{J}$ quantifies the difference between the final SOI feature points at time step $T$ and the targeted SOI state $\mathbf{g}$.
+Within this MPC setup, the cost function $\mathcal{J}$ quantifies the difference between the final SOI observation at time step $T$ and the targeted SOI state $\mathbf{o}^{soi}_{*}$.
 Details on the precise structure of the cost function $\mathcal{J}$ are illustrated in the following section.
 This cost is minimized to yield an optimal sequence of actions across a temporal horizon of $T$ steps:
-$\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{0:{T-1}}$ $=\underset{\langle\mathbf{a}^{0},\mathbf{a}^{1}\rangle_{0:T-1}\in\mathcal{A}}{\arg\min}\mathcal{J}(\mathbf{\hat{g}},\mathbf{g})$ (1)
-$\text{where}\quad\mathbf{\hat{g}}$ $=f_{\theta_{dym}}\left(\mathbf{\hat{o}}_{0},\langle\mathbf{a}^{0},\mathbf{a}^{1}\rangle_{0:{T-1}}\right)$
-where $\hat{\mathbf{g}}$ represents the predicted SOI state at time step $T$. Fig. 4 shows the overall framework of our proposed latent dynamics model for the bimanual bag manipulation task.
-Figure 4: The conceptual representation of the proposed framework for bimanual deformable fabric bag manipulation is based on the latent SOI dynamics model.
-Figure 5: The proposed SOI global particle sampling process.
-Left: the reconstructed point cloud for the fabric bag.
-Right: (a) raw SOI point cloud.
-(b) preprocessed SOI point cloud.
-(c) Reconstructed SOI Surface.
-(d) Resampled SOI particles.
+$\{\mathbf{a}^{0},\mathbf{a}^{1}\}_{0:{T-1}}$ $=\underset{\{\mathbf{a}^{0},\mathbf{a}^{1}\}_{0:T-1}\subseteq\mathcal{A}}{\arg\min}\mathcal{J}(\mathbf{o}^{soi}_{T},\mathbf{o}^{soi}_{*})$ (1)
+$\text{where}\quad\mathbf{o}^{soi}_{T}$ $=f_{\theta_{dyn}}\left(\mathbf{o}^{soi}_{0},\{\mathbf{a}^{0},\mathbf{a}^{1}\}_{0:{T-1}}\right)$
+where $\mathbf{o}^{soi}_{T}$ represents the predicted SOI state at time step $T$.
+Figure 4: The conceptual representation of the proposed framework for bimanual deformable fabric bag manipulation using the latent SOI dynamics model.
+Figure 5: The proposed global particle sampling process: (a) Reconstructed point cloud (PCD) $\mathcal{P}{\text{full}}$; (b) SOI point cloud after HSV filtering $\mathcal{P}{\text{soi}}$; (c) SOI point cloud after down-sampling and outlier removal; (d) Normal generation; (e) Reconstructed SOI surface using ball pivoting; (f) SOI particles after uniform sampling.
 IV Methodology
+In this section, we detail our proposed framework for bimanual deformable bag manipulation using a structure-of-interest (SOI) based neural dynamics model.
+We begin by extracting SOI particles from RGBD data using global particle sampling approach to represent critical bag features, as depicted in Fig.
+4(a).
+A Graph Neural Network (GNN) then models these particles’ dynamics for state prediction (Fig.
+4(b)), which informs the Model Predictive Control (MPC) to adjust robotic gripper actions for shaping the SOI (Fig.
+4(c)).
+This integrated approach streamlines the manipulation process, ensuring target SOI configurations are achieved with precision.
+Input: SOI observation at time step $t$, $\mathbf{o}^{soi}_{t}$, composed of point cloud data $\mathcal{P}^{1}$, $\mathcal{P}^{2}$, $\mathcal{P}^{3}$ and $\mathcal{P}^{4}$ from four different perspectives.
+Output: SOI particle set $\mathbf{P}_{t}$ at time step $t$
+1 for each time step $t$ do
+2       $\mathcal{P}_{full}$leftarrow point-cloud registration based on $\mathbf{o}^{soi}_{t}$ ;
+3       $\mathcal{P}_{soi}$leftarrow extract SOI using $\operatorname{HSV\_filtering}(\mathcal{P}_{full})$;
+4       $\mathcal{P}_{soi}$leftarrow downsampling using $\operatorname{VGD}(\mathcal{P}_{soi},v_{s})$ ;
+5       $\mathcal{P}_{soi}$leftarrow remove outliers using $\operatorname{SOR}(\mathcal{P}_{soi},k)$ ;
+6       $\mathcal{M}_{soi}\leftarrow$ reconstruct surface using $\operatorname{ball\_pivoting}$;
+7       $\mathbf{P}_{t}$leftarrow sample particles using $\operatorname{uniform\_sampling}$ $(\mathcal{M}_{soi})$ ;
+8       return SOI particle set $\mathbf{P}_{t}$ ;
+9
+Algorithm 1 Global Particle Sampling
 IV-A Global Particle Sampling from Raw Observation
-In this work, we begin by sampling representative particles from raw visual inputs, i.e., RGB-D images, to facilitate the training of Graph Neural Network (GNN)-based dynamics model, depicted in Fig.
-4(a).
-The challenge lies in extracting meaningful particle representations of the Structure of Interest (SOI) from visual data that is significantly occluded.
-To this end, we introduce a comprehensive global particle sampling methodology, as illustrated in Fig.
+To address the challenge of real-time extracting meaningful particle representations of the SOI from significantly occluded visual data, we introduce a comprehensive global particle sampling approach, which is illustrated in Fig.
 5.
-Preprocessing: The initial phase of our methodology involves refining the raw point cloud data obtained from RealSense D435 cameras.
-With appropriate calibration, we can transform the raw RGBD images into point cloud data with the cameras’ intrinsic and extrinsic parameters.
-Through meticulous calibration, we can convert raw RGB-D images into point cloud data, utilizing the intrinsic and extrinsic parameters of the cameras.
-To enhance data quality, a statistical outlier removal filter is applied to strip away noise, paving the way for accurate analysis.
-Focusing on the bag’s opening rim, which is crafted from green fabric, we employ a color-based segmentation algorithm.
-A specific color threshold within the HSV color space is set to discern green hues, effectively isolating the rim from the rest of the bag.
-Concurrently, we identify and omit the points marking the bag’s handle areas where robots will interact with the fabric using a separate HSV threshold.
-To ensure the rim’s point cloud is distinctly segregated from other bag components, Euclidean clustering is performed as a concluding step, particularly if handle detection leaves any connected segments.
-This meticulous process results in a purified dataset, optimal for the rim’s geometric representation.
-Surface Reconstruction: To reconstruct the green rim’s surface from its point cloud, we leverage Poisson reconstruction, a preferred technique for its proficiency in managing occlusions by interpolating over incomplete data to form a smooth, continuous surface.
-Traditional methods like ball pivoting and alpha shapes underperform in occluded scenarios.
-Before reconstruction, we compute point normal via local principal axis analysis to guide the Poisson algorithm.
-The resulting surface is further polished with a uniform resampling algorithm, ensuring mesh uniformity and feature preservation for subsequent modeling and manipulation tasks.
-Refinement with Topological Prior: In this step, we ensure that the reconstructed surface conforms to the anticipated topological structure of the bag’s opening rim through a two-fold approach: 1) Topological Analysis: We incorporate topological priors by recognizing that the rim should form a contiguous loop.
-This knowledge is employed to scrutinize the mesh, detecting any topological discrepancies.
-2) Corrective Actions: Upon identifying any topological deviations, we undertake corrective measures by cutting and remeshing to realign the mesh with a loop topology that accurately represents the bag’s rim.
-Resampling: In the final stage of our process, our objective is to generate a high-quality particle set that accurately represents the geometry of the rim.
-To achieve this, we initially implement voxel grid down-sampling to attain a uniform distribution of vertices while concurrently eliminating statistical outliers.
-Subsequently, we employ the farthest point sampling technique to selectively reduce the point cloud associated with the Structure of Interest (SOI) to a manageable quantity conducive to Graph Neural Network (GNN) training.
-The culmination of this process, resulting in a refined particle dataset, is depicted in Fig.
-5(d).
-IV-B SOI Dynamics Model
+Preprocessing: The initial phase of our methodology involves preprocessing of the raw point cloud data obtained from RealSense D435 cameras.
+With appropriate calibration, we can transform the raw RGB-D images into point cloud data (PCD) (see Fig.
+5(a)) with the cameras’ intrinsic and extrinsic parameters.
+Then, we focus on the SOI (i.e., opening rim) of the fabric bag and employ a color-based segmentation algorithm by performing a specific color filtering within the HSV color space to isolate the rim from the rest of the bag (see Fig.
+5(b)).
+To enhance the processing efficiency in the following steps, we further reduce the points to satisfy the real-time performance using a parallel Voxel Grid Downsampling (VGD) by dividing the point cloud into a regular grid of 3D voxels (voxel size denoted by $v_{s}$) and replaces all the points within each voxel with their centroid or average.
+Then, followed by a GPU-accelerated implementation of the Statistical Outlier Removal (SOR) with a $k$-nearest neighbors to remove the noise and further improve the data quality (see Fig.
+5(c)).
+In general, this step is intended to significantly reduce the size and complexity of the point cloud to an appropriate scale (typically, around 500 points) for subsequent surface reconstruction.
+Surface Reconstruction and Refinement: To reconstruct the surface of SOI from its point cloud in a high efficiency, we leverage ball pivoting algorithm that is a preferred technique for its efficiency in managing point clouds.
+To achieve real-time surface reconstruction, we parallelize the algorithm, which allows the BPA to utilize multiple cores or processors simultaneously, thereby accelerating the reconstruction process.
+Specifically, we employ an Octree-based data structure to partition the point cloud into smaller subsets, each of which can be processed independently.
+For example, we can compute point normal (see Fig.
+5(d)) via Octree-based nearest-neighbor search methods in each partitioned subset.
+Then we refine the reconstructed surface based on topological priors to get rid of non-manifold edges and vertices which can be easily detected by Open3D [^1] https://www.open3d.org/docs/release/tutorial/geometry/mesh.html.
+Non-manifold edges and vertices in a mesh are those that do not comply with the manifold criteria, which typically require that each edge belong to exactly two faces and each vertex to form a closed fan of faces around it.
+By removing non-manifold edges and vertices, we can ensure a well-defined volume to further improve the following resampling accuracy.
+Resampling: In the final stage of our process, our objective is to generate a high-quality particle set that accurately represents the geometry of the SOI-based on the reconstructed mesh.
+To achieve this, we employ uniform sampling techniques to selectively reduce the SOI point cloud in to an appropriate fixed scale, maintaining a uniform distribution of the generated particles.
+This ensures a manageable quantity conducive to Graph Neural Network (GNN) training.
+The computation of this process, resulting in a refined particle dataset, is depicted in Fig.
+5(f).
+IV-B SOI-based Neural Dynamics Model
 To characterize the dynamics of SOI particles, as illustrated in Fig.
 4(b), we begin to construct a particle graph and introduce the graph neural networks to model the SOI dynamics.
-To construct a particle graph, we denote a graph representing SOI observations as $\mathcal{G}(\mathbf{\hat{o}}_{t})=(V_{t},E_{t})$, and the graph vertices $V_{t}$ correspond to the SOI’s particles $\mathbf{p}_{i,t}$.
+To construct a particle graph, we denote a graph based the SOI observation to represent the SOI state as $\mathbf{y}_{t}=\mathcal{G}(\mathbf{o}^{soi}_{t})=(\mathbf{P}_{t},\mathbf{E}_{t})$, and the graph vertex set $\mathbf{P}_{t}$ correspond to the SOI’s particles $\mathbf{p}_{i,t}$.
 Each particle is expressed as $\mathbf{p}_{i,t}=\langle\mathbf{x}_{i,t},\mathbf{b}_{i,t}\rangle$, where the position and attributes of the particle $i$ at time $t$ are denoted as $\mathbf{x}_{i,t}$ and $\mathbf{b}_{i,t}$, respectively.
-Edges $E_{t}$ link vertices dynamically, based on spatial relationships, connecting all neighboring particles within a predefined range.
+Attributes of the particles $\mathbf{b}_{i,t}$ are classified into handle particles and non-handle particles.
+The handle particles are identified by a open ball function as below:
+$\mathcal{B}_{r}(\mathbf{x}_{handle})=\{\mathbf{p}_{i}\in\mathbf{P}~{}|~{}d(\mathbf{p}_{i},\mathbf{x}_{handle})<r\}$ (2)
+where $d(\mathbf{p}_{i},\mathbf{x}_{handle})$ represents the distance between SOI particles and bag handle position in Euclidean space.
+The handle position is also identified by HSV filtering in color space.
+Introducing this particle attribute significantly enhances the SOI neural dynamics model.
+The neural dynamics model relies on understanding how different parts (nodes) of the SOI, which lie on the deformable object, behave under robot manipulations.
+By identifying the position of the handle, which has a strong relationship with the enforced robot actions, the model can accurately locate and classify these critical nodes.
+Therefore, node classification allows the model to learn and predict the dynamics of specific particles of the SOI, particularly those near the handle, with greater accuracy.
+Edges $\mathbf{E}_{t}$ link vertices dynamically, based on spatial relationships, connecting all neighboring particles within a predefined range.
 Edge relations are captured by $\mathbf{e}_{j}=\langle m_{j},n_{j},\mathbf{c}_{j}\rangle$, with $1\leq m_{j},n_{j}\leq|\mathbf{P}_{t}|$ denoting the indices of the connected particles, $j$ being the edge index, and $\mathbf{c}_{j}$ describing the type of connection, be it internal structural or handle-to-rim linkages.
-The goal of introducing GNNs is to simulate the dynamics of the SOI and to predict subsequent states from a short historical sequence of SOI particle graphs, formalized as:
-$\mathcal{G}(\mathbf{\hat{o}}_{t+1})=f_{\theta_{\text{dyn}}}\left(\mathcal{G}(\mathbf{\hat{o}}_{t-h:t}),\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{t}\right)$ (2)
+The goal of constructing a forward dynamics model is to predict SOI subsequent state $\mathbf{y}_{t+1}$ based on a short historical sequence of SOI states, denoted by:
+$\mathbf{y}_{t+1}=f_{\theta_{\text{dyn}}}\left(\mathbf{y}_{t-h:t},\mathbf{a}_{t}\right)$ (3)
+With the application of particle graph to represent the SOI states, we can introduce GNNs to simulate the neural dynamics of the SOI to predict subsequent states from a short historical sequence of SOI particle graphs, extended as:
+$\mathcal{G}(\mathbf{o}^{soi}_{t+1})=f_{\theta_{\text{dyn}}}\left(\mathcal{G}(\mathbf{o}^{soi}_{t-h:t}),\{\mathbf{a}^{0},~{}\mathbf{a}^{1}\}_{t}\right)$ (4)
 To facilitate this, the original high-dimensional observation space concerning the SOI is reduced into a condensed, low-dimensional latent graph by encoding the distinct particle and connection features of the SOI.
 The loss function of each encoder is defined based on a distance metric as:
-$\phi^{\mathbf{p}},\psi^{\mathbf{p}}$ $=\arg\min_{\phi^{\mathbf{v}},\psi^{\mathbf{v}}}\operatorname{Dist}[V-(\phi^{\mathbf{p}}\circ\psi^{\mathbf{p}})(P)]$ (3)
-$\phi^{\mathbf{e}},\psi^{\mathbf{e}}$ $=\arg\min_{\phi^{\mathbf{e}},\psi^{\mathbf{e}}}\operatorname{Dist}[E-(\phi^{\mathbf{e}}\circ\psi^{\mathbf{e}})(E)]$
+$\phi^{\mathbf{p}},\psi^{\mathbf{p}}$ $=\arg\min_{\phi^{\mathbf{p}},\psi^{\mathbf{p}}}\operatorname{Dist}[\mathbf{p}_{i}-(\phi^{\mathbf{p}}\circ\psi^{\mathbf{p}})(\mathbf{p}_{i})]$ (5)
+$\phi^{\mathbf{e}},\psi^{\mathbf{e}}$ $=\arg\min_{\phi^{\mathbf{e}},\psi^{\mathbf{e}}}\operatorname{Dist}[\mathbf{e}_{j}-(\phi^{\mathbf{e}}\circ\psi^{\mathbf{e}})(\mathbf{e}_{j})]$
 Here, $\phi^{\mathbf{p}}:V\rightarrow Z_{V}$ and $\phi^{\mathbf{e}}:E\rightarrow Z_{E}$ serve as the encoders for SOI particles and edges, while $\psi^{\mathbf{p}}:Z_{V}\rightarrow V$ and $\psi^{\mathbf{e}}:Z_{E}\rightarrow E$ function as their respective decoders.
@@
 Subsequently, the dynamics are captured using the decoders $\psi^{\mathbf{p}}$ and $\psi^{\mathbf{e}}$, leading to the prediction of the SOI particle graph at time $t+1$:
-$\mathbf{\hat{e}}_{k,t}$ $=\psi^{\mathbf{e}}(z_{j,t}^{\mathbf{e}})_{j=1,\cdots,|E_{t}|}$ (5)
-$\hat{\mathbf{p}}_{i,t+1}$ $=\psi^{\mathbf{p}}\left(z_{i,t}^{\mathbf{p}},\sum_{j\in\mathcal{N}_{i}}\mathbf{\hat{e}}_{k,t}\right)_{i=1,\cdots,|\mathbf{P}_{t}|}$
+$\mathbf{\hat{e}}_{j,t}$ $=\psi^{\mathbf{e}}(z_{j,t}^{\mathbf{e}})_{j=1,\cdots,|\mathbf{E}_{t}|}$ (7)
+$\hat{\mathbf{p}}_{i,t+1}$ $=\psi^{\mathbf{p}}\left(z_{i,t}^{\mathbf{p}},\sum_{j\in\mathcal{N}_{i}}\mathbf{\hat{e}}_{j,t}\right)_{i=1,\cdots,|\mathbf{P}_{t}|}$
 Within this context, $\mathcal{N}_{i}$ denotes the set of edges where particle $i$ is the recipient.
 To adeptly handle the instantaneous propagation of forces, the training also integrates multistep message passing.
-Our training data, derived from point clouds, lacks a consistent point-to-point mapping across each frame, thus precluding the use of particle-wise loss functions.
-To quantify the resemblance between two sets of SOI particle distributions, we investigate two distinct loss functions.
-The first is the widely-adopted Chamfer distance (CD), computed between two particle sets $\mathbf{P}_{1},\mathbf{P}_{2}\subseteq\mathbb{R}^{3}$ as follows:
+Since our training data is based on the particle sets generated from raw point cloud data with our proposed global particle sampling process, it lacks a consistent point-to-point mapping across each frame.
+Therefore, to quantify the similarity between two sets of SOI particle distributions, we investigate two permutation-invariant loss functions.
+The first is the widely adopted Chamfer distance (CD), computed between two particle sets $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ as follows:
 $\small\mathcal{L}_{\text{CD}}(\mathbf{P}_{1},\mathbf{P}_{2})=\sum_{\mathbf{x}_{1}\in\mathbf{P}_{1}}\min_{\mathbf{x}_{2}\in\mathbf{P}_{2}}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{2}^{2}+\sum_{\mathbf{x}_{2}\in\mathbf{P}_{2}}\min_{\mathbf{x}_{1}\in\mathbf{P}_{1}}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{2}^{2}$ (8)
 The second is the Earth mover’s distance (EMD), which is formulated as:
-$\small\mathcal{L}_{\text{EMD}}(\mathbf{P}_{1},\mathbf{P}_{2})=\min_{\phi:\mathbf{P}_{1}\rightarrow\mathbf{P}_{2}}\sum_{\mathbf{x}_{1}\in\mathbf{P}_{1}}\|\mathbf{x}_{1}-\phi(\mathbf{x}_{1})\|_{2}$ (7)
-where $\phi:\mathbf{P}_{1}\rightarrow\mathbf{P}_{2}$ denotes a bijective mapping.
-The EMD represents a solution to the assignment problem, ensuring a unique and stable optimal bijection $\phi$ for almost every pair of particle sets, invariant to infinitesimally small point displacements.
+$\small\mathcal{L}_{\text{EMD}}(\mathbf{P}_{1},\mathbf{P}_{2})=\min_{\lambda:\mathbf{P}_{1}\rightarrow\mathbf{P}_{2}}\sum_{\mathbf{x}_{1}\in\mathbf{P}_{1}}\|\mathbf{x}_{1}-\lambda(\mathbf{x}_{1})\|_{2}$ (9)
+where $\lambda:\mathbf{P}_{1}\rightarrow\mathbf{P}_{2}$ denotes a bijective mapping.
+The EMD represents a solution to the assignment problem, ensuring a unique and stable optimal bijection $\lambda$ for almost every pair of particle sets, invariant to infinitesimally small point displacements.
+Note that the bijection $\lambda$ will be computed between two particle set for every time which means the bijection will be regenerated.
+To ensure real-time performance, we will limit the cardinality of the particle set by configuring the GSP to generate a small number of particles.
+This reduction in particle count helps minimize the computational burden, allowing us to meet the necessary requirements for a real-time controller.
 In essence, EMD in our framework aligns distributions while mitigating point cloud anomalies through the definition of bijective correspondence.
-Regarding the Chamfer distance, it’s worth noting that its use here is somewhat liberal, as it does not fulfill the triangle inequality property.
+Regarding the Chamfer distance, it’s worth noting that its use here is somewhat liberal, as it does not fulfil the triangle inequality property.
 Our composite loss function integrates these distances in a weighted fashion: $\mathcal{L}(\mathbf{P}_{1},\mathbf{P}_{2})=\alpha\mathcal{L}_{\text{CD}}(\mathbf{P}_{1},\mathbf{P}_{2})+\beta\mathcal{L}_{\text{EMD}}(\mathbf{P}_{1},\mathbf{P}_{2})$.
-Empirical evaluations suggest that the optimal weights are $\alpha=0.85$ and $\beta=0.15$.
+Empirical evaluations suggest that the optimal weights are $\alpha=0.15$ and $\beta=0.85$.
 IV-C Model Predictive Control
-Upon training our SOI-centric latent dynamics model, we integrate a model predictive control (MPC) approach to control the robotic gripper in manipulating the fabric bag, as depicted in Fig.
-4(c).
+Figure 6: The MPC framework integrates our trained SOI-based neural dynamics model for bimanual deformable bag manipulation.
+Upon training our SOI-centric neural dynamics model, we integrate a model predictive control (MPC) approach to control the robotic gripper in manipulating the fabric bag, as depicted in Fig.
+10.
 We simplify the gripper’s action space into a parameterized form: $(x,y,z,r_{z})$, with $\{x,y,z\}$ representing the end-effector’s position interacting with the bag handles, and $r_{z}$ indicating the gripper’s rotation around the vertical $z$ axis.
-We omit $r_{x}$ and $r_{y}$ rotations based on empirical evidence suggesting minimal SOI deformation from these movements.
-A goal-oriented MPC is employed as below, using $\mathbf{g}$ to symbolize the desired SOI shape and $\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{0:{T-1}}$ to represent the sequence of action pairs derived via MPC, with $T$ as the planning horizon.
-$\min_{\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{0:T-1}}$ $\mathcal{J}\left(\mathbf{\hat{o}}_{T},\mathbf{g}\right)=\min_{\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{0:T-1}}\mathcal{L}\left(\mathbf{\hat{P}}_{T},\mathbf{P}_{\mathbf{g}}\right)$ (8)
-s.t. $\mathcal{G}(\mathbf{\hat{o}}_{t+1})=f_{\theta_{\text{dym}}}\left(\mathcal{G}(\mathbf{\hat{o}}_{t-h:t}),\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{t}\right)$
-$\mathbf{P}_{\mathbf{g}}=\mathcal{G}(\mathbf{g})$
-$\mathbf{\hat{P}}_{t}=\mathcal{G}(\mathbf{\hat{o}}_{t})$
-$t=0,1,\ldots,T-1$
-The objective is to identify action sequences that minimize the hybrid distance to the goal, expressed as $\mathcal{J}(\mathbf{\hat{g}}_{T},\mathbf{g})\equiv\mathcal{L}(\mathbf{\hat{P}}_{T},\mathbf{P}_{\mathbf{g}})$.
-Gradient-based trajectory optimization is utilized to ascertain the minimal-cost trajectory, showcased in Fig.
-4(c).
-Initial random shooting in the simplified action space is followed by cost calculation via the GNN-driven dynamics model.
-Subsequently, gradient refinement with the limited-memory BFGS method [32] is conducted on the most cost-effective trajectories, employing the same loss function used during the dynamics model’s training phase.
-Figure 6: Experimental set-up.
-(a) Overview of the experimental set-up to validate our SOI-based dynamics model for bimanual deformable bag manipulation tasks.
-(b)-(d) The SOI shape categories are considered in experiments: Long Oval (LO), Round Oval (RO), and Short Oval (SO).
-TABLE I: Averaged SOI particle sampling results across 90 frames.
+We considered only $r_{z}$ in our action space based on our observation that $r_{z}$ induces larger SOI configuration changes compared to $r_{x}$ and $r_{y}$.
+A goal-oriented MPC with $T$ as the planning horizon for the desired SOI observation $\mathbf{o}^{soi}_{*}$, denoted as the particle set $\mathbf{P}_{\mathbf{o}^{soi}_{*}}$, is employed as below:
+$\small\min_{\{\mathbf{a}^{0},~{}\mathbf{a}^{1}\}_{0:T-1}}$ $\sum_{t=0}^{T}\mathcal{L}\left(\mathbf{P}_{\mathbf{o}^{soi}_{t}},\mathbf{P}_{\mathbf{o}^{soi}_{*}}\right)$ (10)
+s.t. $\mathcal{G}(\mathbf{o}^{soi}_{t+1})=f_{\theta_{\text{dyn}}}\left(\mathcal{G}(\mathbf{o}^{soi}_{t-h:t}),\{\mathbf{a}^{0},~{}\mathbf{a}^{1}\}_{t}\right)$
+$d_{lb}\leq||\mathbf{R}^{0}_{t}-\mathbf{R}^{1}_{t}||\leq d_{ub}$
+$||\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{t}||\leq\epsilon$
+where we constrain the distance between left robot end-effector pose $\mathcal{R}^{0}$ and right robot end-effector pose $\mathcal{R}^{1}$ to fall within a specified range from the lower bound $d_{lb}$ to the upper bound $d_{ub}$ so as to prevent potential collisions between the manipulators and safeguard the fabric from tearing due to excessive force.
+Additionally, we ensure that the magnitude of the robot’s actions does not exceed $\epsilon$ to enforce a smooth and continuous trajectory.
+As illustrated in Fig.
+4(c), gradient-based trajectory optimization is utilized to identify the trajectory with the lowest defined cost.
+First, random shooting within the simplified actions is performed, followed by the computation of costs using the SOI-based latent dynamics model.
+Subsequently, we employ the limited-memory BFGS method [28] on the trajectories with the lowest cost to optimize the manipulation actions with gradients, using the same loss function during the dynamics model’s training process.
+The global convergence is guaranteed as proved in [29] and we empirically demonstrated the effectiveness through experiments.
+V Experiments
+V-A Experiment Setup
+As shown in Fig.
+7, we present the general experimental setup used to validate our SOI-based dynamics model for bimanual deformable bag manipulation tasks.
+The setup includes four RealSense D435 RGB-D cameras, each positioned at a corner, to capture RGB-D images of the deformable fabric bag from multiple angles at a frequency of 30Hz and a resolution of 640$\times$480.
+Two Dobot CR5 robotic manipulators, equipped with 7 degrees of freedom, grasp different SOIs using 3D-printed grippers secured with zip ties or binder clips.
+To validate the effectiveness of our proposed approach, three SOIs of the deformable objects, including the opening rim of a bag, the bottom hem and the collar of a T-shirt, have been modified for better perception: it is cut and sewn with the same type of fabric but in contrasting colors, which facilitates the detection while maintaining the consistency of the SOI’s dynamic behavior.
+Additionally, a top-down camera and a front-facing camera are employed to capture the experimental process from various perspectives.
+To support real-time point cloud process, particle sampling and MPC optimization, we utilize a workstation with Ubuntu 20.04, an Intel Core i9-14900K CPU (24 cores, 24 threads), an Nvidia RTX 4090 GPU, 128GB of DDR4/DDR5 RAM, and a 2TB NVMe SSD for storage.
+Table I presents the detailed computation time for the proposed global particle sampling and GNN-based neural dynamics model-integrated MPC optimization process.
+The total computation time required to generate a robot action command is approximately 63.75ms, theoretically allowing the robot to execute up to 15.69 action commands per second to manipulate deformable objects.
+However, to maintain consistency with the action execution frame rate used during the training of our GNN-based neural dynamics model, we standardize the execution frame rate of robot action commands to 10 FPS.
+TABLE I: Computation time of the proposed global particle sampling and MPC optimization process.
+Process Computation Time (ms) $\downarrow$
+Point Cloud Registration $3.13$ $\mathrm{p}\mathrm{m}$  0.26
+SOI Extraction $6.94$ $\mathrm{p}\mathrm{m}$  0.47
+Down-sampling $7.05$ $\mathrm{p}\mathrm{m}$  0.73
+Outlier Removal $8.75$ $\mathrm{p}\mathrm{m}$  0.61
+Ball Pivoting $15.65$ $\mathrm{p}\mathrm{m}$  1.42
+Particle Sampling $6.08$ $\mathrm{p}\mathrm{m}$  0.37
+MPC Computation $17.12$ $\mathrm{p}\mathrm{m}$  1.79
+GNN Inference $0.37$ $\mathrm{p}\mathrm{m}$  0.03
+Figure 7: Experimental set-up to validate bimanual deformable bag manipulation tasks.
+(a) Overview of the set-up.
+(b)-(d) The different SOI shape categories in bag experiments: Long Oval, Round Oval, and Short Oval.
+(e)-(g) The considered bag SOI and two additional types of T-shirt SOIs: opening rim, bottom hem and collar.
+Figure 8: Results of our proposed global particle sampling approach.
+(Top) Reconstructions of different SOIs for different deformable objects.
+(Bottom) Corresponding SOI particle sets.
+(a)-(c): bag opening rim, (d)-(e): T-shirt bottom hem, and (f)-(g): T-shirt collar.
+As illustrated in Fig.
+7(b)-(d), our experiment considers three SOI shape categories for bag manipulation tasks: Long Oval (LO), Round Oval (RO), and Short Oval (SO).
+Additionally, we have incorporated experiments involving two additional types of T-shirt SOIs, each made from distinct fabric materials to further validate the effectiveness of our proposed approach on a variety of deformable objects.
+One SOI is the collar crafted from a soft modal fabric (Fig.
+7(g)), while the other hem SOI comprises a slightly stiffer polyester-cotton blend (Fig.
+7(f)).
+These variations in material properties provide a more comprehensive assessment of our method’s adaptability and robustness across different fabric types.
+For each deformable object, we collect a training dataset of 72,000 frames around 120 minutes, featuring 600 episodes and each spanning 120 frames.
+Within every episode, we execute 20 actions on the deformable objects.
+The data collection policy randomly selects from the action space parameters after doing collision checking and bag constraint checking, which include translations along the $x$, $y$, $z$ axes, and the rotation around $r_{z}$.
+Throughout each episode, we save the generated particle data with proposed sampling approach, as well as the robot end-effector’s poses.
+The experiment comprises two types of manipulation tasks.
+In the SOI shape preservation task, the robotic manipulators aim to keep the SOI shape consistent while moving the fabric bag from its initial shape to a predefined target shape.
+Both the initial and target configurations share the same shape but differ in their spatial transformations.
+Conversely, the SOI shape servoing task involves changing the SOI shape from one configuration to another distinct shape configuration by manipulating the SOIs.
+We conduct a quantitative evaluation for each component of our proposed approach, including SOI particle sampling, the SOI-based neural dynamics model, and performance results of designed manipulation tasks.
+Our primary metrics for evaluating the manipulation performance are Chamfer distance (CD) and Earth mover’s distance (EMD), along with a hybrid metric that combines both.
+Furthermore, we employ Geodesic distance (GD) to assess the proximity of boundary points residing on one manifold with boundary.
+TABLE II: Averaged SOI particle sampling results across 90 frames.
 Sampling Methods CD(cm) $\downarrow$ EMD(cm) $\downarrow$ GD(cm) $\downarrow$
@@
 Local Sampling (LPS) $\mathbf{1.46}\pm 0.11$ $\mathbf{1.08}\pm 0.22$ $\mathbf{1.97}\pm 0.23$
-V Experiments
-V-A Experiment Setup
-As shown in Fig.
-6, we present the general experimental setup used to validate our SOI-based dynamics model for bimanual deformable bag manipulation tasks.
-The setup includes four RealSense D435 RGBD cameras, each positioned at a corner, to capture RGBD images of the deformable fabric bag from multiple angles at a frequency of 30Hz and a resolution of 640$\times$480.
-Two Dobot CR5 robotic manipulators, equipped with 7 degrees of freedom, grasp the bag’s handles using 3D-printed grippers secured with zip ties.
-The Structure of Interest (SOI) of the bag, specifically the opening rim, has been modified for better SOI perception: it is cut and sewn with the same type of fabric but in a contrasting green color, which facilitates the detection while maintaining the consistency of the bag’s dynamic behavior.
-Additionally, a top-down camera and a front-facing camera are employed to capture the experimental process from various perspectives.
-A dataset of 18,000 frames, around 30 minutes, comprises the training data, featuring 200 episodes, each spanning 90 frames.
-Within every episode, we execute six actions on the fabric bag.
-Before each episode’s inception, the fabric bag is orderly arranged to ensure the opening rim protrudes outward without excessive wrinkling.
-The data collection policy indiscriminately selects from the action space parameters after doing collision checking and bag constraint checking, which include translations along the $x$, $y$, $z$ axes, and the rotation around $rz$.
-Throughout each episode, we archive partial point clouds captured by RGB-D cameras, as well as the robot end-effector’s poses as returned by the robot controller.
-Data recording is contingent on the gripper’s movement, an approach taken to optimize memory consumption.
-As illustrated in Fig.
-6(b)-(d), this study considers three SOI shape categories: Long Oval (LO), Round Oval (RO), and Short Oval (SO).
-The experiment comprises two types of manipulation tasks.
-In the SOI shape preservation task, the robotic manipulators aim to keep the SOI shape consistent while moving the fabric bag from its initial position to a predefined target location.
-Both the initial and target configurations share the same shape but differ in their spatial transformations.
-Conversely, the SOI shape servoing task involves altering the SOI shape from one configuration to another distinct shape configuration by manipulating the bag handles.
-We conduct a quantitative evaluation of each component within our framework, encompassing SOI particle sampling, the SOI dynamics model, and performance results of designed manipulation tasks.
-Our primary metrics for evaluating the manipulation performance are Chamfer distance (CD) and Earth mover’s distance (EMD), along with a hybrid metric that combines both.
-Furthermore, we employ Geodesic distance (GD) [20] to assess the proximity of boundary points residing on one manifold with boundary to those on another.
-TABLE II: The performance of the SOI particle dynamics model with different loss function.
-Loss Functions CD(cm) $\downarrow$ EMD(cm) $\downarrow$ GD(cm) $\downarrow$
-CD $1.52\pm 0.19$ $1.94\pm 0.18$ $2.53\pm 0.25$
-EMD $1.85\pm 0.17$ $\mathbf{1.37}\pm 0.15$ $2.76\pm 0.13$
-$0.2$ CD + $0.8$ EMD $1.59\pm 0.16$ $1.40\pm 0.17$ $2.13\pm 0.18$
-$0.1$ CD + $0.9$ EMD $1.55\pm 0.17$ $1.43\pm 0.16$ $2.28\pm 0.19$
-$0.15$ CD + $0.85$ EMD $\mathbf{1.50}\pm 0.16$ $1.38\pm 0.17$ $\mathbf{2.02}\pm 0.16$
-Figure 7: Qualitative results in (a) SOI shape preserving and (b) SOI shape servoing experiments.
 V-B SOI Particle Sampling
 We commence by benchmarking the proposed global particle sampling (GPS) technique against the local particle sampling (LPS) baseline.
-LPS initiates by processing the SOI-specific partial point cloud of the fabric bag through its unique color signature.
-Subsequently, it encapsulates the incomplete SOI-specific cloud with a convex hull.
-Following this, it performs the point sampling procedure, eventually amalgamating the sampled points into a full set of SOI-specific particles.
-As shown in Table I, we compute the mean distance between the sampled particles and the ground-truth particles, the latter captured by a professional 3D scanner.
+LPS initiates by processing the SOI-related partial point cloud of the deformable objects through its color filtering.
+Subsequently, it encapsulates the incomplete SOI point clouds with different convex hulls.
+Following this, it performs the point sampling procedure, eventually combining the sampled points into a full set of SOI particles.
+As shown in Table II, we compute the mean distance between the sampled particles and the ground-truth particles, the latter captured by a professional 3D scanner.
 Our analysis indicates that the GPS method incurs lower losses in terms of all distance metrics, thus outperforming the LPS approach.
-These outcomes align with the premise that incorporating additional topological information about the SOI can markedly enhance the quality of sampling, particularly in scenarios where occlusions are present.
+These outcomes align with the premise that incorporating additional topological information can markedly enhance the quality of sampling, particularly in scenarios where occlusions are present, and Fig.
+8 presents example results for different SOIs applied to three selected deformable objects using our proposed GSP method.
+Figure 9: (Left) SOI-based latent dynamics model training loss with different loss functions in testing data. (Right) The dynamics model’s evaluation over the time horizon in testing data.
+Figure 10: (Left) Mean hybrid distance error and (Right) computation time over different control horizons and gradient steps in the MPC process.
+The white boundary represents the MPC computation time with respect to the horizon and gradient step being less than 20 ms, which is acceptable for the fabric manipulation task.
+Inside the boundary, the computation time for the specific setting of an MPC horizon of 6 and a gradient step of 11 is the most efficient.
 V-C GNN-based SOI Dynamics Model
 The training of our GNN model, detailed in Section IV-B, commences with the construction of a graph, where edges are formed between vertices that are within a distance threshold of $d=0.04$.
+By adopting a 70/15/15 split, we ensure that the GNN model has enough data to learn effectively while also enabling thorough validation and testing.
 Every vertex and edge are encoded using 3-layer Multilayer Perceptrons (MLPs), featuring hidden and output layers, each with 300 neurons.
@@
 These hyperparameters were chosen to foster a robust learning process while maintaining the capacity to capture complex patterns within the data.
-Subsequently, we evaluate the performance of GNN-based SOI dynamics models using different loss functions.
-Table 6 indicates that only using CD or EMD individually optimizes for its own metric but not others, but combining CD and EMD losses leads to better overall performance on all metrics, and the Geodesic distance is significantly improved, compared to using them individually.
-The combination of $0.85~{}\text{CD}+0.15~{}\text{EMD}$ achieves the best performance across all metrics, giving the lowest CD, competitive EMD, and lowest GD, so we select this hybrid distance for the following experiments.
-The reason is that CD and EMD losses capture different aspects of point cloud similarity.
-CD compares point-wise distances while EMD measures global shape differences.
-By combining them, the model is optimized for both local point accuracy and global shape matching.
-The proper combination of 0.85 CD + 0.15 EMD balances these objectives best.
-TABLE III: Mean Hybrid Distance Error for SOI Shape Preserving by Different Dynamics Modeling Methods
+In practice, the final weights for the GNN-based SOI neural dynamics model are chosen based on the lowest validation loss recorded during training.
+This ensures that the model achieves the best possible performance on new, unseen data, which is critical for the practical application of the model in the integrated MPC for the bag manipulation task.
+Subsequently, we evaluate the performance of GNN-based SOI dynamics models using different loss functions in testing data.
+Table III and Fig.
+9(left) indicate that leveraging a hybrid loss function that integrates CD and EMD leads to superior performance across all metrics, as opposed to optimizing individually for either CD or EMD.
+Notably, the Geodesic distance is significantly improved, compared to using them individually.
+The combination of $0.15$ CD + $0.85$ EMD achieves the best performance across all metrics, giving the lowest CD, competitive EMD, and lowest GD, so we select this hybrid distance for the following experiments.
+In Fig.
+9(right), the dynamics model is evaluated on testing data using the CD and the EMD loss over the time horizon.
+Our analysis reveals that while an unavoidable increase in loss metrics occurs as the prediction horizon of the model is extended, these metrics values remain within an acceptable range.
+This outcome validates the potential utility of this model in following deformable object manipulation tasks, demonstrating their capacity to maintain reasonable accuracy over a longer prediction horizon.
+To evaluate the performance of the proposed SOI-based neural dynamics model for real deformable object manipulation tasks, we integrate it with MPC to conduct SOI-preserving experiments on a fabric bag with various oval shapes.
+To balance real-time performance and control accuracy for our manipulation task, we set the MPC horizon to 6 and the gradient step to 11, based on extensive empirical testing (see Fig.
+10).
+This combination provides a reasonable compromise between computational demand and the quality of optimization for the MPC.
+We compare the performance of our proposed approach with two commonly-used and well-established dynamics modeling techniques: the Mass-Spring Model (MSM) [7] and the Finite Element Model (FEM) [8].
+These two molding techniques are also integrated into the MPC framework for conducting the experiments.
+The comparative analysis over 60 trails includes both quantitative and qualitative outcomes, as shown in Table IV, Fig.
+11(a) and Fig.
+12(a).
+Our findings reveal a clear trend across all examined methods; there is an incremental rise in error corresponding to the transformation of the bag’s shape from a Long Oval (LO) to a Short Oval (SO), suggesting that as the bag opening rim becomes shorter and wider, the complexity of preserving its structure increases.
+Notably, our GNN-based dynamics model consistently outperforms the MSM and FEM across all object shapes, achieving the lowest hybrid loss in tasks dedicated to shape preservation.
+This outcome indicates the superior expressive capability of our GNN-based model in capturing the intricate deformations of the fabric bags, and maintaining the SOI shape unchanged during the fabric bag moving, surpassing the traditional MSM and FEM approaches, particularly when dealing with objects that exhibit more complex dynamic behaviors.
+The advantage of our model is most pronounced when interacting with simpler geometric shapes.
+TABLE III: The performance of the SOI particle dynamics model with different loss functions in testing data
+Loss Functions CD(cm) $\downarrow$ EMD(cm) $\downarrow$ GD(cm) $\downarrow$
+CD $1.52\pm 0.19$ $1.94\pm 0.18$ $2.53\pm 0.25$
+EMD $1.85\pm 0.17$ $\mathbf{1.37}\pm 0.15$ $2.76\pm 0.13$
+$0.2$ CD + $0.8$ EMD $1.59\pm 0.16$ $1.40\pm 0.17$ $2.13\pm 0.18$
+$0.1$ CD + $0.9$ EMD $1.55\pm 0.17$ $1.43\pm 0.16$ $2.28\pm 0.19$
+$0.15$ CD + $0.85$ EMD $\mathbf{1.50}\pm 0.16$ $1.38\pm 0.17$ $\mathbf{2.02}\pm 0.16$
+TABLE IV: Mean Hybrid Distance Error for SOI Shape Preserving by Different Dynamics Modeling Methods over 60 Trials
 Method Long Oval (LO) Round Oval (RO) Short Oval (SO)
@@
 Ours $\textbf{1.69}\pm 0.18$ $\textbf{1.84}\pm 0.22$ $\textbf{2.08}\pm 0.21$
-TABLE IV: Mean Hybrid Distance Error and Success Rate for SOI Shape Servoing by Different Methods
-Method Mean Hybrid Distance Error Success Rate
-LO $\rightarrow$ SO SO $\rightarrow$ RO LO $\rightarrow$ RO LO $\rightarrow$ SO SO $\rightarrow$ RO LO $\rightarrow$ RO Total
-VS [33] $3.75\pm 0.19$ $3.69\pm 0.22$ $3.92\pm 0.23$ $21/30$ $19/30$ $22/30$ $68.89\%$
-LSC [34] $2.94\pm 0.48$ $3.14\pm 0.42$ $2.89\pm 0.44$ $24/30$ $24/30$ $26/30$ $82.23\%$
-MSM [35] $3.22\pm 0.43$ $3.49\pm 0.47$ $3.28\pm 0.39$ $19/30$ $21/30$ $23/30$ $70.00\%$
-FEM [36] $2.61\pm 0.36$ $2.82\pm 0.39$ $2.67\pm 0.33$ $23/30$ $22/30$ $26/30$ $78.89\%$
-Ours $2.25\pm 0.23$ $2.39\pm 0.28$ $2.13\pm 0.24$ $28/30$ $/30$ 30 $29/30$ 96.67%
+TABLE V: Mean Hybrid Distance Error and Success Rate for SOI Shape Preserving on Different Deformable Objects by Different Methods
+Method Bag Opening Rim (180 trls) Soft T-shirt Collar (60 trls) Stiff T-shirt Bottom Hem (60 trls) Total
+Mean HD Error $\downarrow$ Success Rate $\uparrow$ Mean HD Error $\downarrow$ Success Rate $\uparrow$ Mean HD Error $\downarrow$ Success Rate $\uparrow$
+VS [30] $3.20\pm 0.20$ $71.11\%$ $3.41\pm 0.21$ $76.67\%$ $3.28\pm 0.29$ $68.33\%$ $71.67\%$
+MSM [7] $2.80\pm 0.37$ $73.89\%$ $2.48\pm 0.40$ $71.67\%$ $3.06\pm 0.48$ $65.00\%$ $71.67\%$
+FEM [8] $2.17\pm 0.34$ $83.33\%$ $2.23\pm 0.33$ $80.00\%$ $2.64\pm 0.38$ $86.67\%$ $83.33\%$
+G-DOOM [24] $3.07\pm 0.32$ $79.44\%$ $2.64\pm 0.34$ $75.00\%$ $3.31\pm 0.43$ $71.67\%$ $77.00\%$
+Ours $1.86\pm 0.23$ $\mathbf{98.33\%}$ $1.65\pm 0.27$ $\mathbf{95.00\%}$ $2.43\pm 0.26$ $\mathbf{86.67\%}$ $\mathbf{95.33\%}$
+TABLE VI: Mean Hybrid Distance Error and Success Rate for SOI Shape Servoing on Different Deformable Objects by Different Methods
+Method Bag Opening Rim (180 trls) Soft T-shirt Collar (60 trls) Stiff T-shirt Bottom Hem (60 trls) Total
+Mean HD Error $\downarrow$ Success Rate $\uparrow$ Mean HD Error $\downarrow$ Success Rate $\uparrow$ Mean HD Error $\downarrow$ Success Rate $\uparrow$
+VS [30] $3.76\pm 0.21$ $68.89\%$ $3.75\pm 0.19$ $71.67\%$ $3.84\pm 0.27$ $70.00\%$ $69.67\%$
+MSM [7] $3.29\pm 0.45$ $70.00\%$ $2.94\pm 0.42$ $66.67\%$ $3.57\pm 0.47$ $63.33\%$ $68.00\%$
+FEM [8] $2.70\pm 0.35$ $78.89\%$ $2.62\pm 0.37$ $76.67\%$ $2.88\pm 0.39$ $80.00\%$ $78.67\%$
+G-DOOM [24] $3.48\pm 0.33$ $75.56\%$ $3.11\pm 0.37$ $68.33\%$ $3.87\pm 0.41$ $66.67\%$ $72.33\%$
+Ours $2.27\pm 0.24$ $\mathbf{96.67\%}$ $1.87\pm 0.26$ $\mathbf{91.67\%}$ $2.68\pm 0.29$ $\mathbf{90.00\%}$ $\mathbf{94.33\%}$
+Figure 11: Qualitative results in (a) SOI shape preserving and (b) SOI shape servoing tasks with our proposed approach.
+Figure 12: Achieved SOI particle configurations of (a) SOI shape preserving and (b) SOI shape servoing tasks by different approaches.
 V-D Manipulation Results
-To evaluate the performance of the proposed SOI-based dynamics model for deformable fabric bag manipulation, we conducted SOI preserving and servoing experiments on a fabric bag with different oval shapes and the transitions between them.
-In SOI preserving experiments, we mainly compare the performance of our proposed approach with two commonly-used and well-established dynamics modeling techniques: the Mass-Spring Model (MSM) [35] and the Finite Element Model (FEM) [36].
-The comparative analysis encompasses both quantitative and qualitative outcomes, as encapsulated in Figure 7(a) and Table III.
-Our findings reveal a discernible trend across all examined methods; there is an incremental rise in error corresponding to the transformation of the bag’s shape from a Long Oval (LO) to a Short Oval (SO), suggesting that as the bag opening rim becomes shorter and wider, the complexity of preserving its structure increases.
-Notably, our GNN-based dynamics model consistently outperforms the MSM and FEM across the entire spectrum of object shapes, achieving the lowest error rates in tasks dedicated to shape preservation.
-This outcome corroborates the superior expressive capability of our GNN-based model in capturing the intricate deformations of the fabric bags, and maintaining the SOI shape unchanged during the fabric bag moving, surpassing the traditional MSM and FEM approaches, particularly when dealing with objects that exhibit more complex dynamic behaviors.
-The advantage of our model is most pronounced when interacting with simpler geometric shapes.
-In our SOI servoing experiments, we conduct a comprehensive analysis of the manipulation results associated with transitions between distinct shape categories of the structure of interest (SOI).
-Additionally, we juxtapose our dynamics model with two sophisticated shape servoing techniques tailored for deformable objects–visual servoing (VS) as described by Lagneau et al.
-[33], and the latent shape control (LSC) model introduced by Qi et al.
-[34].
-The analysis integrates both quantitative and qualitative assessments, as depicted in Fig.
-7(b) and detailed in Table IV.
-We consider three specific shape servoing tasks: LO (Long Oval) $\rightarrow$ SO (Short Oval), SO $\rightarrow$ RO (Round Oval), and LO $\rightarrow$ RO.
-The evaluation metrics include the Mean Hybrid Distance Error (recorded solely for successful trials) and the Success Rate of SOI servoing.
-Furthermore, we report the success rate for achieving complete servoing with a threshold error of less than $5$ cm, based on 30 trials for each method and shape transition scenario.
-The experimental outcomes reveal that our proposed method consistently secured the least shape error, ranging from $2.13$ to $2.39$ cm across all tested cases, and achieved an outstanding overall success rate of $96.67\%$.
-When contrasted with traditional model-based techniques such as the Mass-Spring Model (MSM) and the Finite Element Model (FEM), our data-driven SOI dynamics model exhibited a markedly enhanced capability in manipulating the SOI for the fabric bag.
-It was observed that transformations involving shorter and wider shapes incurred higher errors due to their inherently complex dynamics.
-Nevertheless, our proposed method demonstrated a robust performance across all tasks.
-In conclusion, the bimanual manipulation experiments underscore the effectiveness of the proposed SOI-based dynamics modeling approach, which ensures precise and dependable control over the shapes of deformable objects through advanced predictive modeling and optimization techniques.
+[…]
 VI Conclusion
-In this work, we introduced a novel bimanual manipulation framework for deformable bags, centered around a Structure of Interest (SOI)-based latent dynamics model.
+In this work, we introduced a novel bimanual manipulation framework for deformable bags, centered around a SOI-based latent dynamics model.
 Our approach effectively integrates multi-view perception, graph neural networks, and model predictive control, facilitating precise and efficient robotic manipulation of flexible materials.
-Through simulations and real-world experiments, the framework demonstrated promising results in intelligent physical interaction with deformable objects.
+Through real-world experiments, the framework demonstrated promising results in intelligent physical interaction with deformable objects.
 One limitation of the current system is its reliance on differently colored SOIs to distinguish target manipulation areas, which may not be feasible in all operational settings.
-Future efforts will aim to enhance the system’s ability to identify and manipulate SOIs without such visual aids, broadening the applicability of our method to a wider array of real-world scenarios.
+Future efforts will focus on employing a learning-based SOI detection module and multimodal sensing to enhance the system’s ability to identify and manipulate SOIs without such visual aids, thereby broadening the applicability of our method to a wider array of real-world scenarios.